### v1

# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

Abstract
The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts. In this paper, we study the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm over directed graphs (digraphs). Unlike methods that rely solely on row-stochastic or column-stochastic mixing matrices, Push-Pull avoids nonlinear correction and has shown superior empirical performance across a variety of settings. However, its theoretical analysis remains challenging, and the linear speedup property has not been generally established—revealing a significant gap between empirical success and limited theoretical understanding. To bridge this gap, we propose a novel analysis framework and prove that Push-Pull achieves linear speedup over arbitrary strongly connected digraphs. Our results provide the comprehensive theoretical understanding for stochastic Push-Pull, aligning its theory with empirical performance. Code: https://github.com/pkumelon/PushPull.
1 Introduction
Scaling machine learning tasks to large datasets and models necessitates efficient distributed computation across multiple nodes. In distributed optimization, a network of $n$ nodes cooperates to solve
$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F(x;\xi_{i})].$ (1)
Here, $\xi_{i}$ denotes a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function, which may be nonconvex. During optimization, each node $i$ can access its own local gradient $\nabla F_{i}(x;\xi_{i})$, but must communicate to obtain gradient information from the other nodes. Data heterogeneity is typically present in (1), meaning the local distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ differ across nodes. Consequently, the loss functions $f_{i}(x)$ and $f_{j}(x)$ are generally not equal, i.e., $f_{i}(x)\neq f_{j}(x)$.
Parallel stochastic gradient descent (SGD) is a widely adopted approach in distributed optimization, where local gradients are globally averaged using either a central server [1] or Ring All-Reduce communication primitives [2]. However, such global synchronization incurs substantial bandwidth overhead and high latency as the number of nodes increases, thereby hurting the scalability of the optimization process. To address these limitations, decentralized SGD has recently been proposed to solve problem (1) without requiring global gradient aggregation. In this setting, each node independently maintains and updates its model by communicating only with immediate neighbors, significantly reducing communication overhead. Beyond efficiency, decentralized algorithms offer improved robustness to node and link failures, maintaining convergence as long as the underlying communication graph remains connected. In contrast, centralized SGD breaks down if the central server crashes.
In this paper, we consider a general decentralized setting where the network topology is modeled as a strongly connected directed graph (digraph), which naturally encompasses undirected networks as a special case. Compared to undirected graphs, digraphs are particularly well-suited for ad hoc decentralized systems, such as heterogeneous robotic networks that operate under asynchronous, one-sided communication and random node activation. More recently, digraphs have shown great potential in distributed deep learning, where certain classes of directed networks exhibit both topological sparsity and strong algebraic connectivity. These properties make them especially attractive for the distributed training of large-scale models. Notable examples include the exponential graph [3], EquiTopo [4], and B-ary tree [5].
1.1 Decentralized algorithms over digraphs
In any connected network, the topology can be characterized by a mixing matrix. This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance. For undirected networks, it is straightforward to construct symmetric and doubly stochastic mixing matrices. For directed networks, achieving a doubly stochastic matrix is generally infeasible. Consequently, mixing matrices in such cases are typically either column-stochastic [6, 7] or row-stochastic [8, 9], but not both.
Building on the connectivity pattern of the underlying network topology, decentralized optimization algorithms over directed graphs can be broadly categorized into three families in the literature. The first family relies on the Push-Sum technique [10, 11], which employs column-stochastic mixing matrices for all communications. Accordingly, we refer to this class as Push-Only algorithms in this paper. When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [12, 11] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity. More advanced methods such as EXTRA-Push [13], DEXTRA [14], ADD-OPT [15], and Push-DIGing [7] improve convergence speed by mitigating the effects of data heterogeneity. Stochastic variants, including SGD-Push [16] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting. Just as Push-Sum underpins the design of Col-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods. Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting. Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23]. In this framework, a row-stochastic matrix is used to exchange model parameters, while a column-stochastic matrix is employed to share gradient information. As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices. [25] extended the Push-Pull method to stochastic optimization settings, while [5] improved its convergence rate on B-ary tree network topologies. Accelerated or constrained variant of the Push-Pull method have also been explored in [26, 27].
1.2 Linear speedup in iteration complexity
Linear speedup refers to the property of a distributed algorithm whereby the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes. For example, centralized SGD achieves an $\epsilon$-accurate solution in $\mathcal{O}\left(\textstyle\frac{1}{n\epsilon^{2}}\right)$ iterations , where $n$ denotes the number of nodes. This scaling behavior demonstrates that centralized SGD achieves linear speedup with respect to the number of nodes. The linear speedup property is crucial for establishing the advantage of distributed algorithms over their single-node counterparts. Without this property, the practical utility of a distributed algorithm may be limited, as it would offer little benefit from using multiple computing nodes.
The linear speedup property of decentralized algorithms over undirected graphs is well established. The first such analysis for decentralized SGD was presented in [28], demonstrating that decentralized SGD can achieve the same asymptotic convergence rate as its centralized counterpart. Building on this foundation, a variety of advanced decentralized algorithms—such as Exact Diffusion/NIDS [29, 30] and Gradient Tracking methods [31, 32, 33]—are proven to achieve linear speedup [34, 29, 35]. In addition, [36] provides a linear speedup analysis for decentralized SGD under time-varying undirected network topologies.
Recent literature has increasingly focused on establishing the linear speedup property of decentralized algorithms over digraphs. For Push-Only algorithms that employ column-stochastic mixing matrices, linear speedup has been formally proven in [16, 37]. Moreover, a recent study [38] proposes methods that achieve the optimal iteration complexity in this setting. For Pull-Only algorithms that employ row-stochastic mixing matrices, convergence has been extensively studied under various settings [9, 18, 19, 20, 21]; however, these analyses do not establish the linear speedup property. The main challenge lies in the fact that using a row-stochastic matrix alone introduces a deviation in the descent direction from the globally averaged gradient. With a novel analysis framework addressing this issue, the linear speedup property of Pull-Only algorithms was established only recently in [39]. These results reveal the effectiveness of Push-Only and Pull-Only algorithms over digraphs.
However, although Push-Pull algorithms typically achieve significantly better empirical performance than Push-Only and Pull-Only methods, their linear speedup property has not been generally established. Existing theoretical analyses of Push-Pull remain limited, primarily focusing on strongly convex loss functions in deterministic settings [23, 40, 22, 5]. In the nonconvex and stochastic setting, neither linear speedup nor even basic convergence guarantees have been formally established for Push-Pull. The most relevant theoretical result to date is presented in [5], which introduces a variant called B-ary Tree Push-Pull (BTPP). While BTPP is shown to achieve linear speedup, the result relies on the specialized B-ary tree topology, and the analysis does not generalize to general network topologies.
1.3 Motivation and main results
Building on the above discussion, the existing theoretical convergence analyses for Push-Pull methods remain rather limited, underscoring a clear discrepancy between the weak theoretical guarantees and their strong empirical performance. This gap gives rise to the following two open questions:
  • Q1.
    What is the intrinsic difference between the Push-Pull method and the other two families, i.e., Push-Only and Pull-Only? Why is the theoretical analysis of Push-Pull more challenging?
  • Q2.
    Can Push-Pull achieve linear speedup over a broader class of network topologies—for instance, all strongly connected digraphs?
Addressing these two questions is crucial for establishing a theoretical foundation that supports the effectiveness of Push-Pull methods.
Regarding Question Q1, we demonstrate that both Push-Only and Pull-Only algorithms closely align with centralized SGD; specifically, after a careful numerical transformation, their variable updates closely resemble those of centralized SGD. This structural similarity explains why the linear speedup property can be readily established for Push-Only and Pull-Only methods. In contrast, Push-Pull is not designed to align with centralized SGD and consistently exhibits a non-vanishing deviation from the centralized trajectory. Consequently, the linear speedup analysis for Push-Pull is substantially more challenging and fundamentally distinct from that of Push-Only and Pull-Only algorithms.
For Question Q2, we establish a linear speedup convergence result for the Push-Pull method. To the best of our knowledge, this is the first such result for Push-Pull over all strongly connected digraphs under standard assumptions, whereas existing analyses of stochastic Push-Pull [5, 41] are restricted to tree-structured topologies. By introducing a novel analysis framework that addresses the misalignment between Push-Pull and centralized SGD, this paper bridges the gap between the convergence theory and the strong empirical performance of the Push-Pull method.
The remainder of the paper is organized as follows. Section 2 introduces necessary notations, assumptions, and measures. Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods, and offers insights into the derivation of linear speedup convergence. In Section 4, we present the key components of our analysis and state the main convergence theorem. Section 5 provides extensive numerical experiments that demonstrate the ability of the Push-Pull method to achieve linear speedup. Finally, Section 6 concludes the paper.
2 Notations, Assumptions and Measures
This section introduces notations, assumptions, and measures that will be used throughout the paper.
Notations. The directed communication network is denotes as ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, where ${\mathcal{N}}=\{1,2,\ldots,n\}$ is the set of nodes, and ${\mathcal{E}}$ is the set of directed edges. An edge $(j,i)\in{\mathcal{E}}$ means that node $i$ can send information to node $i$. We say a directed graph is strongly connected if there is a directed path from any node to all other nodes in the graph. We say that ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is the underlying topology of weight matrix $A$ if $A$ is entrywise nonnegative, and $a_{ij}>0$ if and only if $(j,i)\in{\mathcal{E}}$.
Let $\mathds{1}\in\mathbb{R}^{n}$ denote the vector of all-ones. The set $[n]$ denotes the indices $\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ signifies the diagonal matrix whose diagonal elements are comprised of $v$.
We use matrix $A$ to denote the row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$). The Perron vectors [42] for matrices $A$ and $B$, if they exist, are denoted as $\pi_{A}$ and $\pi_{B}$, respectively. We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
$\displaystyle{\mathbf{x}}^{(k)}$ $\displaystyle:=[({x}_{1}^{(k)})^{\top};({x}_{2}^{(k)})^{\top};\cdots;({x}_{n}^{(k)})^{\top}]$
$\displaystyle{\mathbf{g}}^{(k)}$ $\displaystyle:=[\nabla F_{1}({x}_{1}^{(k)};\xi_{1}^{(k)})^{\top};\cdots;\nabla F_{n}({x}_{n}^{(k)};\xi_{n}^{(k)})^{\top}]$
$\displaystyle\nabla f({\mathbf{x}}^{(k)})$ $\displaystyle:=[\nabla f_{1}({x}^{(k)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(k)}_{n})^{\top}]$
[…]
Assumptions. We first specify the underlying graphs (${\mathcal{G}}_{A}$, ${\mathcal{G}}_{B}$) and weight matrices ($A$, $B$) used in stochastic Push-Pull.
Assumption 1 (Communication Graph).
Graphs ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ are strongly connected.
Assumption 2 (Weight Matrix).
Weight matrix $A$ is row-stochastic, with an underlying topology of ${\mathcal{G}}_{A}$. Weight matrix $B$ is column-stochastic, with an underlying topology of ${\mathcal{G}}_{B}$. The traces of $A$ and $B$ are positive.
Assumptions 1 and 2 are commonly used in the analysis of Push-Pull [24, 40, 26]. In our paper, the primary role of these assumptions is to derive Proposition 1. Strictly speaking, replacing these assumptions while retaining Proposition 1 would not affect our main convergence results.
Proposition 1 (Perron Vector, Exponential Decay).
For $A$ and $B$ satisfying Assumption 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
$\pi_{A}^{\top}A=\pi_{A}^{\top},\quad B\pi_{B}=\pi_{B},\quad\|\pi_{A}\|_{1}=\|\pi_{B}\|_{1}=1,\quad\pi_{A}^{\top}\pi_{B}>0.$
Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ satisfying
$\|A^{k}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{k},\quad\|B^{k}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{k},\quad\forall k\geq 0.$
Proof.
The proof follows directly from the Perron-Frobenius theorem [42]. ∎
Finally, we state standard assumptions regarding the stochastic gradients and function smoothness.
Assumption 3 (Gradient Oracle).
There exists a constant $\sigma\geq 0$ such that
$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)}),\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2},\quad\forall i\in[n],\,k\geq 0.$
Assumption 4 (smoothness).
There exist positive constants $L,\Delta$ such that
$\displaystyle\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|\quad\mbox{and}\quad f_{i}(x)-\inf_{x}f_{i}(x)\leq\Delta,\quad\forall i\in[n].$
Measures. When Proposition 1 holds, we can define the following constants
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\quad s_{B}:=\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\quad c:=n\pi_{A}^{\top}\pi_{B}.$ (2)
In this paper, we use only $s_{A}$, $s_{B}$, $c$, and $n$ to characterize the influence of the mixing matrices. Existing analyses on directed graphs typically employ measures related to the induced matrix norm [25, 38, 39]. In contrast, our measures only require the $\ell_{2}$ norm. An interesting relationship between our measures and other measures is presented in Proposition 2.
Proposition 2 (Relation with other measures).
Suppose Assumption 1 and 2 hold. We have the following estimate on $s_{A}$, $s_{B}$ and $c$:
$\displaystyle 1\leq s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{1+\frac{1}{2}\ln(\kappa_{B})}{1-\beta_{B}},$
$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\},$
where $\beta_{A}:=\|\Pi_{A}^{\frac{1}{2}}(A-A_{\infty})\Pi_{A}^{-\frac{1}{2}}\|_{2}\in[0,1)$, $\beta_{B}:=\|\Pi_{B}^{-\frac{1}{2}}(B-B_{\infty})\Pi_{B}^{\frac{1}{2}}\|_{2}\in[0,1),$ $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$, $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ are measures from literature [25, 38, 39], $\Pi_{A}=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}=\mathrm{Diag}(\pi_{B})$.
Proof.
See Appendix A. ∎
3 Preliminary
In this section, we provide a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms. Furthermore, we discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
Push-Only and Pull-Only Algorithms Align with Centralized SGD. Most decentralized algorithms, such as decentralized SGD [36], decentralized gradient tracking [43, 14], (sub)gradient-Push [44], Push-Diging [7], (sub)gradient-Pull [9], and FROST [19], are designed to mimic centralized SGD. These algorithms can generally be formulated as:
$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-\alpha\bar{g}^{(k)}+\,\epsilon^{(k)}$ (3)
where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}x_{i}^{(k)}$ is a convex combination of local decision variables, and $\bar{g}^{(k)}=\frac{1}{n}\sum_{i=1}^{n}\nabla F_{i}(x_{i}^{(k)},\xi_{i}^{(k)})$ represents the average gradient. Here, $\epsilon^{(k)}$ denotes an additional error term.
For Push-Only and Pull-Only optimization algorithms, the extra error term is elaborately designed to be zero or to approach zero exponentially fast. Such design ensures that the trajectories of decision variables closely resemble those of centralized SGD. Given the sufficiently small extra error, the linear speedup result naturally follows within the same analytical framework used for centralized SGD. The following is a brief example of how to obtain the formulation (3).
Example 1.
Push-Diging [7]. The algorithm can be written as:
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A_{k}{\mathbf{x}}^{(k)}-\alpha V_{k+1}^{-1}{\mathbf{y}}^{(k)},$ (4a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (4b)
where $V_{k}=\mathrm{Diag}(v_{k})=\mathrm{Diag}(B^{k}\mathds{1})$ is a diagonal matrix and $A_{k}=V_{k+1}^{-1}BV_{k}$. By left-multiplying $n^{-1}v_{k+1}^{\top}$ on both sides of (4a) and omitting some calculation details, we will obtain
$n^{-1}v_{k+1}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}.$
By taking $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k+1)}$, this is exactly the form of (3).
Figure 1: Distributed logistic regression on synthetic dataset. We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005. The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant. For detailed settings, see 36, Appendix C.
Push-Pull Analysis Suffers from Extra Error Terms. Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (5a)
$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (5b)
[…]
$\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$
where $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$ is the projected decision variable, $c=n\pi_{A}^{\top}\pi_{B}$ is a constant. We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\bar{g}^{(k)}$ is the dominating factor. But here, the error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|$ introduces additional noise that may dominate the noise in $\bar{g}^{(k)}$.
For certain specific topologies, the noise in the error term can either be ignored or absorbed. For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull and demonstrates the linear speedup result on this particular topology. In their setting, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, which leads to the expression $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$. By setting the loss function at the first node to zero, they obtain the relationship $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$, which simplifies the process of achieving linear speedup.
Unfortunately, for general topologies, the error term lacks exploitable structure, making direct analysis overly conservative. This leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result. Consequently, the central challenge becomes: How can we accurately characterize the noise within this error term? We tackle this key question in the following section.
4 Achieving Linear Speedup for Push-Pull
For simplicity, we denote $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$. Using the gradient tracking recursion (5b), we discover the following relationship about extra error terms: for any integer $k\geq 0,m\geq 1$:
$\displaystyle\sum_{i=1}^{m}\Delta_{y}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}).$ (6)
Equation (6) reveals a contraction-type recursion governing the sum of consecutive error terms. Informally, if we restrict our analysis to terms involving $\sigma^{2}$, we observe that $\|{\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}\|F^{2}={\mathcal{O}}(\sigma^{2})$. This contraction property further implies that the aggregated term $\|\sum{i=1}^{m}\Delta_{y}^{(k+i)}\|_{F}^{2}$ maintains the same ${\mathcal{O}}(\sigma^{2})$ scaling. Consequently, we can deduce that individual terms satisfy $\|\Delta_{y}^{(k)}\|_{F}^{2}={\mathcal{O}}(\frac{\sigma^{2}}{m})$.
This scaling relationship becomes particularly significant when $m$ is sufficiently large: in this regime, the $\sigma^{2}$-dependent component of the extra error term becomes negligible. As a result, our analysis preserves the crucial linear speedup property. To formalize this logic, we start from the following contraction relationship.
Lemma 3.
For any integers $k\geq 0$ and $m\geq 1$, we have
$\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (7)
Furthermore, we can define a ”noiseless” version for the right-hand side, i.e.
$H_{m,k}:=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),$
and the distance therein can be bounded by:
$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\varepsilon^{(k+i)}-H_{m,k}\|_{F}^{2}]\leq 2ns_{B}^{2}\sigma^{2}.$ (8)
The proof of Lemma 3 is provided in Appendix B.1. This lemma implies that the summation of the $m$ error terms can be represented as $H_{m,k}$ plus a bounded noise term. Moreover, the term $H_{m,k}$ exclusively involves difference of deterministic gradients and a single instance of ${\mathbf{y}}^{(k)}$, making it straightforward to estimate its magnitude.
To effectively leverage the result of Lemma 3, we introduce an $m$-step descent lemma. This new lemma follows by an estimation on $H_{m,k}$ and applying the standard $L$-smoothness descent lemma to the following recursion:
$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}.$ (9)
Lemma 4 (m-step Descent Lemma).
Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. The following inequality holds for any integers $k\geq 0,m\geq 1$:
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (10)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+\frac{L}{16m}\Delta_{1}^{(k)}+\frac{L}{16m}\Delta_{2}^{(k)}+{3\alpha^{2}Ls_{B}^{2}}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}],$
where
$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
Remarkably, if we divide both sides of (10) by $\frac{c\alpha m}{4}$, the first noise term simplifies to $\frac{8\alpha cL}{n}\sigma^{2}$, which corresponds to the standard linear speedup term. The second noise term simplifies to $\frac{8\alpha ns_{B}^{2}}{cm}\sigma^{2}$. When we select some $m$ larger than $c^{-2}n^{2}s_{B}^{2}$, the second noise term should have a smaller order than the first term, thereby ensuring the achievement of the standard linear speedup, which verifies our logic before. Next, we try to estimate $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$.
Lemma 5 (Consensus Error).
Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{3c}{16mL},\frac{1}{6s_{B}L}\}$, we have
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}.$
where $\Delta_{x}:={\mathbf{x}}^{(k)}-A_{\infty}{\mathbf{x}}^{(k)}$, $\Delta_{y}^{(k)}:={\mathbf{y}}^{(k)}-B_{\infty}{\mathbf{y}}^{(k)}$.
Lemma 5 transfers the complicated error terms to commonly-seen consensus error terms. Now, we can apply techniques similar to [38] to bound $\Delta_{x}$ and $\Delta_{y}$.
Lemma 6.
Suppose Assumption 34 and Proposition 1 hold. For any $T\geq 1$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2},$
and
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}].$
Now we are ready to state the main result of this paper.
Theorem 1.
Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$ and $T:=mK$. Let $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$. When the learning rate $\alpha$ satisfies $\alpha\leq\min\{\frac{1}{8cL},\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L}\}$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$
where $C_{1}$ is a constant determined by $A$ and $B$. Furthermore, if we choose
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\},$
we will obtain that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{C_{\Delta}L\Delta}{T},$ (11)
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.
Comparison with existing analyses. Theorem 1 provides the first finite-time non-convex convergence result for Push-Pull, while previous analyses focus either on the strongly convex case [45, 40] or on variants of Push-Pull relying on specific topologies [5]. The first term $\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary strongly connected topologies.
Impact of $m$. In our analysis, we consider a $m$-step gradient descent. The number $m$ effectively controls the $\sigma^{2}$-related term in the extra error. A common question is can we choose an arbitrarily large $m$ in this analysis. We can do that, but choosing a larger $m$ will force us to select a smaller learning rate, which finally leads to a larger transient time. To sum up, number $m$ provides a chance for tradeoff between the noise level and the transient time.
Transient time. Our theorem provides a coarse estimate for the transient time, which is of ${\mathcal{O}}(nC_{\Delta}^{2})$. We do not recommend using this to estimate the real transient time, but it still reflects the impact of topology. For example, if we use a directed ring graph where $s_{A}=s_{B}={\mathcal{O}}(n)$, $c={\mathcal{O}}(1)$, the transient time can be ${\mathcal{O}}(n^{17})$. If we use a directed exponential graph [3] where $s_{A}=s_{B}={\mathcal{O}}(1)$, $c={\mathcal{O}}(1)$, the transient time will be ${\mathcal{O}}(n^{5})$.
5 Experiment
In this section, we empirically investigate the linear speedup properties of the Push-Pull algorithm across diverse network topologies. For the stochastic gradient oracle, we ensure that each node operates on a local dataset, computing gradients from randomly sampled data points at each iteration. The experimental validation proceeds in two phases: first, we examine a nonconvex regression problem using synthetic data, followed by training a fully connected neural network on the heterogeneously distributed MNIST dataset.
To further assess practical performance on established benchmarks, we extend our evaluation to ResNet-18 training on CIFAR-10. The distributed training experiments are conducted across multiple nodes ($n>1$) using three distinct network topologies: directed exponential graphs, grid graphs, and undirected random graphs. All results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup achieved through distributed optimization.
Complete implementation specifications are provided in Appendix C. The source code and experimental scripts are publicly available in our GitHub repository.
5.1 Nonconvex Logistic Regression on Synthetic Dataset
In the first experiment, the loss function at each node is a logistic regression function with nonconvex regularization terms (see (36) in Appendix C). We studied six different network topologies: the exponential, grid, and ring graphs are directed, while the random, geometric, and nearest neighbor graphs are undirected. The mixing matrices $A$ and $B$ are generated by a special procedure (see Appendix C.1), which produces row-stochastic or column-stochastic matrices, but not doubly-stochastic ones.
We can observe clear linear speedup curves in Figure 2. For each topology, the noise level in gradients decreases proportionally to the square root of the number of nodes after the same number of communication rounds. All learning rates are set to $lr=0.005$.
Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all 0.0050.005 0.0050.005.
5.2 Training Neural Network for MNIST Classification
In the second experiment, we trained a three-layer fully connected neural network for MNIST classification. We used the same topologies as before: directed exponential, grid, and ring graphs, and undirected random, geometric, and nearest neighbor graphs. Data is distributed heterogeneously using a Directlet distribution, see Appendix C.3. Figure 3 shows that Push-Pull is able to achieve linear speedup on both directed and undirected topologies.
Figure 3: Linear speedup performance over various topologies, the learning rates are the same over the same topology.
5.3 Training Neural Network for CIFAR10 Image Classification
In the third experiment, we trained Resnet18 for CIFAR10 classification across different numbers of nodes and various topologies. We used three topologies mentioned before: directed exponential, grid graphs, and undirected random graphs. Here $n=1$ means training a single Resnet18 on a single-node without distributed optimization.
Figure 4 demonstrates the performance of our Push-Pull method. We observe that Push-Pull’s accuracy is initially lower compared to single-node optimization, which can be attributed to the overhead of the influence of topology. Nevertheless, after a sufficient number of epochs, Push-Pull achieves superior accuracy, leveraging its inherent speed advantages.
Figure 4: The test accuracy across different numbers of nodes and various topologies. All experiments were conducted using a learning rate of 0.0050.005 0.0050.005. The lower three figures specifically depict the accuracy after 100100 100100 epochs.
6 Conclusion
In this paper, we introduce a novel analysis technique and demonstrate linear speedup convergence for the Push-Pull method across arbitrary strongly connected topologies. This discovery enhances the community’s comprehension of the Push-Pull method and validates its effectiveness. Numerical experiments have corroborated our theoretical findings.
Appendix A Rolling-Sum Lemma
Lemma 7 (Rolling Sum Lemma).
For a row-stochastic weight matrix satisfying Proposition 1, the following inequality holds for any $T\geq 0$:
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}^{2}\sum_{i=0}^{T}\|\Delta^{(i)}\|_{F}^{2},$ (12)
where $\Delta^{(i)}\in\mathbb{R}^{n\times d}$ are arbitrary matrices, $s_{A}$ is defined as in (2). Furthermore, if Assumption 1 and 2 hold, we have
$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$
where $\beta_{A}$, $\kappa_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
Proof.
First, we prove that
$\displaystyle\|A^{i}-A_{\infty}\|_{2}\leq\sqrt{\kappa_{A}}\beta_{A}^{i},\forall i\geq 0.$ (13)
Notice that $\beta_{A}:=\|A-A_{\infty}\|_{\pi_{A}}$ and
$\|A^{i}-A_{\infty}\|_{\pi_{A}}=\|(A-A_{\infty})^{i}\|_{\pi_{A}}\leq\|A-A_{\infty}\|^{i}_{\pi_{A}}=\beta_{A}^{i},$
we have
$\|(A^{i}-A_{\infty})v\|=\|\Pi_{A}^{-1/2}(A^{i}-A_{\infty})v\|_{\pi_{A}}\leq\sqrt{\underline{\pi_{A}}}\beta_{A}^{i}\|v\|_{\pi_{A}}\leq\sqrt{\kappa_{A}}\beta_{A}^{i}\|v\|,$
where $\underline{\pi_{A}}:=\min\pi_{A}$, $\overline{\pi_{A}}:=\max\pi_{A}$. The last inequality comes from $\|v\|_{\pi_{A}}\leq\overline{\pi_{A}}\|v\|$. Therefore, (13) holds.
Second, we want to prove that
$\displaystyle s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$ (14)
Towards this end, we define $M_{A}:=\max\{1,\max_{k\geq 0}\|A^{k}-A_{\infty}\|_{2}\}$. According to (13), $M_{A}$ is well-defined. We also define $p=\max\left\{\frac{\ln(\sqrt{\kappa_{A}})-\ln(M_{A})}{-\ln(\beta_{A})},0\right\}$, then we can verify that $\|A^{i}-A_{\infty}\|_{2}\leq\min\{M_{A},M_{A}\beta_{A}^{i-p}\},\forall i\geq 0.$ With this inequality, we can bound $\sum_{i=0}^{k}\|A^{k+1-i}-A_{\infty}\|_{2}$ as follows:
$\displaystyle\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}=\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}\|A^{i}-A_{\infty}\|_{2}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}\|A^{i}-A_{\infty}\|_{2}$ (15)
[…]
If $p=0$, (15) is simplified to $\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}\cdot\frac{1}{1-\beta_{A}}$ and (14) is naturally satisfied. If $p>0$, let $x=\min\{\lfloor p\rfloor,k\}+1-p\in[0,1)$, (14) is simplified to
$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}}).$
By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This finishes the proof of (14).
Finally, to obtain (12), we use Jensen’s inequality. For any positive numbers $a_{0},a_{1},\ldots,a_{k}$ satisfying $\sum_{i=0}^{k}a_{i}=1$, we have
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}=\|\sum_{i=0}^{k}a_{k-i}\cdot a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}$ (16)
[…]
By choosing $a_{k-i}=(\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2})^{-1}\|A^{k-i}-A_{\infty}\|_{2}$ in (16), we obtain that
$\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\cdot\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}.$ (17)
By summing up (17) from $k=0$ to $T$, we obtain that
$\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}\sum_{k=0}^{T}\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}$
[…]
which finishes the proof of this lemma. ∎
Appendix B Proofs of Main Theorem
B.1 Proof of Lemma 3
For $1\displaystyle 1\leq i\leq m-$ 1, consider
$\displaystyle(I-B_{\infty}){\mathbf{y}}^{(k+i)}$ $\displaystyle=(I-B_{\infty})(B{\mathbf{y}}^{(k+i-1)}+{\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k+i-1)})$
$\displaystyle=(B-B_{\infty}){\mathbf{y}}^{(k+i-1)}+(I-B_{\infty}){\mathbf{g}}^{(k+i)}-(I-B_{\infty}){\mathbf{g}}^{(k+i-1)}$
Sum it up for $1\displaystyle 1\leq i\leq m-$ 1, we obtain that
$(I-B_{\infty})\sum_{i=1}^{m-1}{\mathbf{y}}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-2}{\mathbf{y}}^{(k+i)}+(I-B_{\infty}){\mathbf{g}}^{(k+m-1)}-(I-B_{\infty}){\mathbf{g}}^{(k)}.$
Similarly, consider $(B-B_{\infty}){\mathbf{y}}^{(k+i)}$, $2\displaystyle 1\leq i\leq m-$ 2, we have
$(B-B_{\infty})\sum_{i=1}^{m-2}{\mathbf{y}}^{(k+i)}=(B^{2}-B_{\infty})\sum_{i=0}^{m-3}{\mathbf{y}}^{(k+i)}+(B-B_{\infty}){\mathbf{g}}^{(k+m-2)}-(B-B_{\infty}){\mathbf{g}}^{(k)}.$
Repeat this procedure for $m-3,m-4,\ldots,1$ and sum them up, we finally arrive at
$\sum_{i=0}^{m-1}(I-B_{\infty}){\mathbf{y}}^{(k+i)}=\sum_{i=0}^{m-1}(B^{i}-B_{\infty}){\mathbf{y}}^{(k)}+\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
which is just the first statement. Next, note that the difference between $\sum_{i=0}^{m-1}\varepsilon^{(k+i)}$ and $H_{m,k}$ equals to:
$\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))-\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})).$
By taking the squared norm, we obtain that:
$\displaystyle\quad\sum_{i=1}^{m-1}\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}$
$\displaystyle+\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}+\text{Inner Product Terms}.$
When we take the expectation conditioned on ${\mathcal{F}}_{k}$, all the inner product terms become zero. For example, for any $1\displaystyle 0\leq i<j\leq m-$ 1,
$\displaystyle\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\right\rangle\right]$
$\displaystyle=\mathbb{E}_{k}[\mathbb{E}_{k+j}[\text{This inner product}]]$
$\displaystyle=\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),0\right\rangle\right]=\mathbb{E}[0]=0.$
Therefore, we only need to estimate the squared terms. By the Cauchy-Schwarz inequality, and $\|\pi_{A}\|\leq 1$, we have
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|_{F}^{2}]$
$\displaystyle\leq\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}\leq ns_{B}^{2}\sigma^{2},$
and
$\displaystyle\quad\mathbb{E}_{k}[\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$
$\displaystyle\leq\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq\|\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}n\sigma^{2}\leq ns_{B}^{2}\sigma^{2}.$
By combining the two parts, we finish the proof. $\square$
B.2 Proof of Lemma 4
For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x^{(k)}_{i})$. According to Assumption 3, we have
$\mathbb{E}[\bar{g}^{(k+i)}]=\mathbb{E}[\mathbb{E}_{k}[\bar{g}^{(k+i)}]]=\mathbb{E}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
Note that
$\pi_{A}^{\top}\sum_{i=0}^{m-1}{\mathbf{y}}^{(k+i)}=c\sum_{i=0}^{m-1}\bar{g}^{(k+i)}+\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$
the descent lemma on $\hat{x}^{(k+m)}=\hat{x}^{(k)}-\alpha\pi_{A}^{\top}\sum_{i=k}^{k+m-1}{\mathbf{y}}^{(i)}$ can be written as follows:
$\displaystyle f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})$ $\displaystyle\leq-c\alpha\left\langle\sum_{i=0}^{m-1}\bar{g}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle-\alpha\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$
$\displaystyle\quad+c^{2}\alpha^{2}L\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}+\alpha^{2}L\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}.$
By taking expectation (from time 00 0), we obtain that
$\displaystyle\mathbb{E}[f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})]\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (18)
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
$\displaystyle\quad+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
The first inner product term in (18) can be bounded by:
$\displaystyle-\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle$ $\displaystyle=-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{2}\|\overline{\nabla f}^{(k+i)}-\nabla f(\hat{x}^{(k)})\|^{2}$
$\displaystyle\leq-\frac{1}{2}\|\overline{\nabla f}^{(k+i)}\|^{2}-\frac{1}{2}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{L^{2}}{2}\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness. The first quartic term in (18) can be bounded by:
$\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
where we use the fact that $\bar{g}^{(k+i)}$ are conditional unbiased estimate and
$\mathbb{E}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
Plug these estimations into (18), we obtain that
$\displaystyle\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (19)
$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
Next, we try to simplify the last quadratic term in (19) . Notice that
$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}),$
the quadratic term can be bounded as:
$\displaystyle\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (20)
$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
The first term here can be estimated using the fact that $\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\leq s_{B}^{2}$. For the second term in (20), according to Assumption 3, we have
$\displaystyle\quad\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (21)
$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq n\sigma^{2}\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+n\sigma^{2}\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}^{2}$
$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
The inner product can produce a similar term. According to Lemma 3,
$\mathbb{E}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}[H_{m}({\mathbf{y}}^{(k)},\nabla f({\mathbf{x}}^{(k)},\ldots,\nabla f({\mathbf{x}}^{(k+m-1)}))].$
Therefore, we have
$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (22)
$\displaystyle=-\alpha\mathbb{E}\left[\left\langle\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\quad+\mathbb{E}\left[\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right){\mathbf{y}}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$
$\displaystyle\leq\frac{1}{cmL}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\quad+\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}].$
Plug (20) and (22) into (19), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (23)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
For $2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
$\displaystyle\quad\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}].$ (24)
Take (24) into (23), we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (25)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
We can further apply Jensen inequality on the last term in (25), i.e.
$\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
$\displaystyle\leq(\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2})^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
$\displaystyle\leq s_{B}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
Therefore, (25) can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
By taking $\alpha\leq\min\{\frac{1}{8cmL},\frac{c}{8n^{2}s_{B}^{2}L}\}$, this can be further simplified to
$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (26)
$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{L}{16m}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
$\displaystyle+\frac{L}{16m}\sum_{j=1}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
which finishes the proof.
B.3 Proof of Lemma 5
We consider how to measure the difference of ${\mathbf{x}}$. Use the fact that
$\displaystyle{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
We decompose ${\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ and apply the Cauchy-Schwarz inequality:
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]\leq 3\left(\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\right)\|\Delta_{x}^{(k)}\|_{F}^{2}$ (27)
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$
$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}+3\alpha^{2}s_{B}^{2}L^{2}\Delta_{2}^{(k)},$
where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness.
For ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, similarly, we decompose it into
$\displaystyle{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})+\Delta_{x}^{(k)},$
which finally produces
$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]\leq 4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (28)
$\displaystyle\quad+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4\alpha^{2}s_{B}^{2}L^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]+4m\|\Delta_{x}^{(k)}\|_{F}^{2}.$
Inequalities (27) and (28) tell us
$\displaystyle\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}$
where $v_{1}=3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$, $v_{2}=4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4m\|\Delta_{x}^{(k)}\|_{F}^{2}$ are positive numbers. To proceed on, we assume $\alpha\leq\frac{1}{6s_{B}L}$, so that the inversion of the $2\times 2$ coefficient matrix satisfies
$\displaystyle\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\leq{\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}}^{-1}\leq\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}.$
Therefore, we have
$\displaystyle{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{10m}\\ \frac{L}{10m}\end{pmatrix}}^{\top}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}.$
This gives that
$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2},$
which finishes the proof of Lemma 5.
B.4 Proof of Lemma 6
In this section, we provide some inequalities for the consensus error $\Delta_{x}$ and $\Delta_{y}$. These inequalities will only be used in the final part of proof of our main theorem.
Lemma 8 (Consensus of x).
Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, we have
$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]$
$\displaystyle\quad+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2}$
Proof.
First note that
$\displaystyle\Delta_{x}^{(k+1)}$ $\displaystyle=(I-A_{\infty}){\mathbf{x}}^{(k+1)}=(A-A_{\infty}){\mathbf{x}}^{(k)}-(I-A_{\infty}){\mathbf{y}}^{(k)}$
$\displaystyle=\ldots=\sum_{i=0}^{k}(A^{k-i}-A_{\infty}){\mathbf{y}}^{(i)}$
$\displaystyle=\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta_{y}^{(i)}+\sum_{i=0}^{k}(A^{k-i}-A_{\infty})n\pi_{B}\bar{g}^{(i)}$
Using the rolling sum lemma, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{x}^{(k+1)}\|_{F}^{2}\leq 2s_{A}^{2}\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}+2n^{2}s_{A}^{2}\|\pi_{B}\|^{2}\sum_{k=0}^{T}\|\bar{g}^{(k)}\|_{F}^{2}.$
Note that $\mathbb{E}[\|\bar{g}^{(k)}\|_{F}^{2}]\leq n^{-1}\sigma^{2}+\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$ and $\|\pi_{B}\|\leq 1$, we finish the proof. ∎
Lemma 9 (Consensus of y).
Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
Proof.
$\displaystyle\Delta_{y}^{(k+1)}$ $\displaystyle=(B-B_{\infty})\Delta_{y}^{(k)}+(I-B_{\infty})({\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)})$
$\displaystyle=\ldots=(B^{k+1}-B_{\infty}){\mathbf{g}}^{(0)}+\sum_{i=0}^{k}(B^{k-i}-B_{\infty})({\mathbf{g}}^{(i+1)}-{\mathbf{g}}^{(i)})$
Using the rolling sum lemma, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq s_{B}^{2}\|{\mathbf{g}}^{(0)}\|+s_{B}^{2}\sum_{k=0}^{T}\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}$
By Cauchy-Schwarz Inequality and our assumption for bounded noise, we have
$\displaystyle\mathbb{E}[\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 6n\sigma^{2}+3\mathbb{E}[\|\nabla f({\mathbf{x}}^{(k+1)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 6n\sigma^{2}+3L^{2}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]$
Besides, $\mathbb{E}[\|{\mathbf{g}}^{(0)}\|_{F}^{2}]\leq n\sigma^{2}+\|\nabla f({\mathbf{x}}^{(0)})\|_{F}^{2}\leq n\sigma^{2}+2nL\Delta$. Therefore, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7)\sigma^{2}+2ns_{B}^{2}L\Delta+3s_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$ (29)
Since
$\displaystyle\quad\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=\mathbb{E}[\|(A-I){\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)}\|_{F}^{2}]$ (30)
$\displaystyle=\mathbb{E}[\|(A-I)\Delta_{x}^{(k)}-\alpha\Delta_{y}^{(k)}+n\alpha\pi_{B}\bar{g}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq 3\|A-I\|_{2}^{2}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+3\alpha^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+3n\alpha^{2}\sigma^{2}+3n^{2}\alpha^{2}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}],$
and $\|A-I\|_{2}^{2}\leq 2n$, we can apply (30) to substitute the last term in (29). This produces
$\displaystyle(1-9\alpha^{2}L^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T)\sigma^{2}$
$\displaystyle\quad+9ns_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+9\alpha^{2}L^{2}n^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]+2ns_{B}^{2}L\Delta$
Using the conclusion of Lemma 8, this can be further transformed to
$\displaystyle\quad(1-9\alpha^{2}L^{2}s_{B}^{2}-18\alpha^{2}L^{2}ns_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$
$\displaystyle\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T+18\alpha^{2}L^{2}ns_{A}^{2}T)\sigma^{2}+2ns_{B}^{2}L\Delta$
$\displaystyle\quad+(9\alpha^{2}L^{2}n^{2}s_{B}^{2}+18\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
When $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, this can be simplified to
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
B.5 Proof of Theorem 1
Plug Lemma 5 into Lemma 4, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (31)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\left(\frac{2\alpha^{2}c^{2}mL}{n}+4\alpha^{2}Lns_{B}^{2}\right)\sigma^{2}$
$\displaystyle\quad+\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}.$
When $\alpha\leq\frac{1}{8cmL}$, $\frac{\alpha c(1-4\alpha mcL)}{2}$ can be replaced with $\frac{c\alpha}{4}$. Sum up (31) for $k\in[m^{*}]$, take full expectation and divide both sides by $\frac{c\alpha T}{4}$ we obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (32)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{4\alpha Lns_{B}^{2}}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{4s_{A}^{2}L}{cm\alpha T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+\frac{16\alpha Ls_{B}^{2}}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{x}^{(k+1)}\|_{F}^{2}$, we can plug in Lemma 8 into (32), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (33)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{16\alpha L(s_{A}^{4}+s_{B}^{2})}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Similarly, since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$, we further plug in Lemma 9 into (33), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (34)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+C_{\sigma}\sigma^{2}+\frac{128ns_{B}^{2}\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}+C_{f}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
where
$\displaystyle C_{\sigma}$ $\displaystyle=\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}+\frac{480\alpha Lns_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cm},$
$\displaystyle C_{f}$ $\displaystyle=\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}+\frac{2000\alpha^{3}L^{3}n^{3}s_{A}^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cmT}.$
When $m\geq 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$, we have $C_{\sigma}\leq\frac{30c\alpha L}{n}$ and $C_{f}\leq\frac{2c\alpha L}{5T}+\frac{100c\alpha^{3}L^{3}n^{3}s_{A}^{2}}{T}$. When we further set $\alpha\leq\min\{\frac{1}{8cL},\frac{1}{4ns_{A}L}\}$, we have $C_{f}\leq\frac{1}{T}$. Under these conditions, (34) can be simplified to
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$ (35)
Finally, to satisfy all the constraints on $m$ and $\alpha$, we can set $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$, and
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\}.$
Plugging this into (35) gives that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+C_{\Delta}\frac{L\Delta}{T},$
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.
Appendix C Experiment Details
In this section, we supplement experiment setup details.
C.1 Weight Matrix Design
We use the following procedure to generate a row/column-stochastic weight matrix. For a fixed number of nodes $n$, we begin by initializing a 0-1 adjacency matrix $W$ as the identity matrix $I$ to account for self-loops on each node. Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$. After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from 11 11 to 99 99 to all entries where $W_{i,j}=1$. Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
C.2 Synthetic data
For experiments conducted on the synthetic dataset, the objective function is given by:
$\displaystyle f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad x\in\mathbb{R}^{d},$ (36)
where
$f_{i}(x)=\underbrace{\frac{1}{L_{i}}\sum_{l=1}^{L_{i}}\ln\left(1+\exp(-y_{i,l}h_{i,l}^{\top}x)\right)}_{\text{Logistic Loss}}+\underbrace{\rho\sum_{j=1}^{d}\frac{[x]_{j}^{2}}{1+[x]_{j}^{2}}}_{\text{Non-convex Regularization}},\quad L_{i}=L_{\text{Local}}/n.$
Data Synthesis Process:
  1. 1.
    Global Data Generation:
    • Generate optimal parameters: $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$
    • Create global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with $h_{l,j}\sim\mathcal{N}(0,1)$
    • Compute labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ via randomized thresholding:
      $y_{l}=\begin{cases}1&\text{if }1/z_{l}>1+\exp(-h_{l}^{\top}x_{\text{opt}})\\ -1&\text{otherwise}\end{cases},\quad z_{l}\sim\mathcal{U}(0,1)$
  2. 2.
    Data Distribution:
    • Partition $H$ and $L$ across $n$ nodes using:
      $H^{(i)}=H[iM:(i+1)M,\;:\;],\;\quad Y^{(i)}=Y[iM:(i+1)M],\;\quad M=L_{\text{total}}/n$
      where $L_{\text{total}}$ must be divisible by $n$
  3. 3.
    Local Initial Points:
    $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d}),\quad\sigma_{h}=10$
Gradient Computation: At each iteration, each node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$. The gradient computation consists of two components:
$\nabla f_{i,B}(x)=\underbrace{-\frac{1}{B}\sum_{b=1}^{B}\frac{y_{i,b}h_{i,b}}{1+\exp(y_{i,b}h_{i,b}^{\top}x)}}_{\text{Logistic Loss Gradient}}+\underbrace{\rho\sum_{j=1}^{d}\frac{2[x]_{j}}{(1+[x]_{j}^{2})^{2}}}_{\text{Non-convex Regularization}}$
where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly from the $L_{i}$ local samples without replacement. Notice that the gradient on each node $i$ is computed on the local parameter $x_{i}\in\mathbb{R}^{d}$.
Implementation Details:
  • Global dataset size $L_{\text{total}}=204800$, batch size $B=200$
  • Dimension $d=10$, regularization $\rho=0.01$
  • Node configurations: $n\in\{4,6,8,9,12,16,18,24,25\}$, and the least common multiple (LCM) is $3,600$.
Notice that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
Evaluation Metric: We track $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ as gradient norm, where
$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\;M=L_{\text{Total}}/n.$
The experimental results shown in our plots represent the average performance across 2020 2020 independent repetitions, where each trial is conducted with the fixed random $\text{seed}=42$ for reproducibility.
C.3 Neural Network Experiment
We first employ three-layer fully connected neural networks for MNIST classification.
Data Distribution: For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector
$\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$
where $\mathbf{1}^{10}$ is a 1010 1010-dimensional vector of ones. Each $q_{k,c}$ represents the target proportion of samples from class $c$ that node $k$ should ideally hold. By definition of the Dirichlet distribution, $q_{k,c}\geq 0,\forall c$ and $\sum_{c=0}^{9}q_{k,c}=1,\forall k$.
Denote $N_{k,c}^{\text{target}}=q_{k,c}\cdot\frac{M}{n}$ as the target number of samples of class $c$ to be assigned to node $k$. Ignoring the slight deviations caused by non-divisibility, the number of samples assigned to each node $k$ is approximately $60,000/n$. Here we set $\alpha=0.9$ to get the heterogeneous data and the Figure 5 shows the distribution of samples with 1616 1616 nodes. The Figure 2 shows that the Push Pull achieves linear speedup on synthetic data.
Figure 5: Comparison of uniform (left) and heterogeneous (right) data distributions across 1616 1616 nodes ($\alpha=0.9$).
Evaluation Metric: After each batch computation, we first average the gradients obtained from the $n$ neural networks. Then we compute the norm of this averaged gradient. This norm is subsequently normalized by the number of parameters P of a single neural network. In our experiments, $\textbf{P}=1,863,690$.
$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{ij}}\right\|_{2}/\sqrt{\textbf{P}},\text{ the gradient }\mathbf{g}_{i,B_{i}}\text{ is computed on the $\displaystyle j$-th batch of node $\displaystyle i$}$ (37)
The Figure 3 shows that the Push Pull achieves linear speedup on both directed and undirected topologies.
Finally, we trained Resnet18 models on the CIFAR10 dataset to evaluate the performance of our Push-Pull method on a benchmark computer vision task. This involved conducting training across different numbers of nodes ($n>1$) utilizing the three previously described network topologies: directed exponential, grid graphs, and undirected random graphs. The performance was compared to single-node training ($n=1$) without distributed optimization. (see Figure 4).
Implementation Details:
  • Learning rate $lr=0.005$ on each topology and each number of node.
  • Batch size $bs=128$ on each topology and each number of node.

### v2

@@
 The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting.
-Just as Push-Sum underpins the design of Col-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods.
+Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods.
 Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting.
@@
 Notations.
-The directed communication network is denotes as ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$, where ${\mathcal{N}}=\{1,2,\ldots,n\}$ is the set of nodes, and ${\mathcal{E}}$ is the set of directed edges.
-An edge $(j,i)\in{\mathcal{E}}$ means that node $i$ can send information to node $i$.
-We say a directed graph is strongly connected if there is a directed path from any node to all other nodes in the graph.
-We say that ${\mathcal{G}}=({\mathcal{N}},{\mathcal{E}})$ is the underlying topology of weight matrix $A$ if $A$ is entrywise nonnegative, and $a_{ij}>0$ if and only if $(j,i)\in{\mathcal{E}}$.
-Let $\mathds{1}\in\mathbb{R}^{n}$ denote the vector of all-ones.
-The set $[n]$ denotes the indices $\{1,2,\ldots,n\}$.
-For a vector $v$, $\mathrm{Diag}(v)$ signifies the diagonal matrix whose diagonal elements are comprised of $v$.
-We use matrix $A$ to denote the row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$).
-The Perron vectors [42] for matrices $A$ and $B$, if they exist, are denoted as $\pi_{A}$ and $\pi_{B}$, respectively.
+The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ is the set of directed edges.
+An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$.
+A directed graph is said to be strongly connected if there exists a directed path from any node to every other node in the graph.
+We say that $\mathcal{G}=(\mathcal{N},\mathcal{E})$ is the underlying topology of the weight matrix $A=[a_{ij}]$ if $A$ is entrywise nonnegative and $a_{ij}>0$ if and only if $(j,i)\in\mathcal{E}$.
+Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector.
+Set $[n]$ denotes $\{1,2,\ldots,n\}$.
+For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ as its diagonal entries.
+We use matrix $A$ to denote the row-stochastic matrix (i.e., $A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix (i.e., $B^{\top}\mathds{1}=\mathds{1}$).
+The Perron vectors [42] of matrices $A$ and $B$, if they exist, are denoted by $\pi_{A}$ and $\pi_{B}$, respectively.
 We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$.
@@
 […]
-Assumptions. We first specify the underlying graphs (${\mathcal{G}}_{A}$, ${\mathcal{G}}_{B}$) and weight matrices ($A$, $B$) used in stochastic Push-Pull.
+Assumptions. We first specify the underlying graphs ($\mathcal{G}_{A}$, $\mathcal{G}_{B}$) and the corresponding weight matrices ($A$, $B$) used in the stochastic Push-Pull method.
 Assumption 1 (Communication Graph).
@@
 Assumption 2 (Weight Matrix).
-Weight matrix $A$ is row-stochastic, with an underlying topology of ${\mathcal{G}}_{A}$.
-Weight matrix $B$ is column-stochastic, with an underlying topology of ${\mathcal{G}}_{B}$.
-The traces of $A$ and $B$ are positive.
-Assumptions 1 and 2 are commonly used in the analysis of Push-Pull [24, 40, 26].
-In our paper, the primary role of these assumptions is to derive Proposition 1.
-Strictly speaking, replacing these assumptions while retaining Proposition 1 would not affect our main convergence results.
-Proposition 1 (Perron Vector, Exponential Decay).
-For $A$ and $B$ satisfying Assumption 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
+The weight matrix $A$ is row-stochastic with underlying digraph $\mathcal{G}_{A}$, and the weight matrix $B$ is column-stochastic with underlying digraph $\mathcal{G}_{B}$.
+Both $A$ and $B$ have positive trace.
+Assumptions 1 and 2 are standard in the analysis of the Push-Pull method [24, 40, 26].
+In our paper, their primary role is to facilitate the derivation of Proposition 1.
+Strictly speaking, these assumptions could be replaced as long as Proposition 1 remains valid, without affecting the main convergence results.
+Proposition 1 (Exponential Decay).
+For $A$ and $B$ satisfying Assumptions 1 and 2, there exist entrywise positive Perron vectors $\pi_{A},\pi_{B}\in\mathbb{R}^{n}$ such that
 $\pi_{A}^{\top}A=\pi_{A}^{\top},\quad B\pi_{B}=\pi_{B},\quad\|\pi_{A}\|_{1}=\|\pi_{B}\|_{1}=1,\quad\pi_{A}^{\top}\pi_{B}>0.$
-Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ satisfying
+Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
 $\|A^{k}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{k},\quad\|B^{k}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{k},\quad\forall k\geq 0.$
@@
 There exists a constant $\sigma\geq 0$ such that
-$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)}),\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2},\quad\forall i\in[n],\,k\geq 0.$
+$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)})\quad\mbox{and}\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2}$
+for any $i\in[n]$ and $k\geq 0$.
 Assumption 4 (smoothness).
@@
 $\displaystyle\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|\quad\mbox{and}\quad f_{i}(x)-\inf_{x}f_{i}(x)\leq\Delta,\quad\forall i\in[n].$
-Measures. When Proposition 1 holds, we can define the following constants
-$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\quad s_{B}:=\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\quad c:=n\pi_{A}^{\top}\pi_{B}.$ (2)
-In this paper, we use only $s_{A}$, $s_{B}$, $c$, and $n$ to characterize the influence of the mixing matrices.
-Existing analyses on directed graphs typically employ measures related to the induced matrix norm [25, 38, 39].
-In contrast, our measures only require the $\ell_{2}$ norm.
-An interesting relationship between our measures and other measures is presented in Proposition 2.
+Measures.
+Various effective metrics have been proposed to capture the impact of digraphs on decentralized stochastic optimization.
+Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$.
+Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{1/2}(B-B_{\infty})\Pi_{B}^{-1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$.
+Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
+However, this paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define the following constants:
+$\displaystyle s_{A}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\},$ (2)
+$\displaystyle s_{B}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\},$ (3)
+$\displaystyle c$ $\displaystyle:=n\,\pi_{A}^{\top}\pi_{B}.$ (4)
+Throughout this paper, we characterize the influence of the mixing matrices using only $s_{A}$, $s_{B}$, $c$, and $n$.
+A connection between these quantities and the previously discussed metrics is established in Proposition 2.
 Proposition 2 (Relation with other measures).
-Suppose Assumption 1 and 2 hold. We have the following estimate on $s_{A}$, $s_{B}$ and $c$:
-$\displaystyle 1\leq s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{1+\frac{1}{2}\ln(\kappa_{B})}{1-\beta_{B}},$
-$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\},$
-where $\beta_{A}:=\|\Pi_{A}^{\frac{1}{2}}(A-A_{\infty})\Pi_{A}^{-\frac{1}{2}}\|_{2}\in[0,1)$, $\beta_{B}:=\|\Pi_{B}^{-\frac{1}{2}}(B-B_{\infty})\Pi_{B}^{\frac{1}{2}}\|_{2}\in[0,1),$ $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$, $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ are measures from literature [25, 38, 39], $\Pi_{A}=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}=\mathrm{Diag}(\pi_{B})$.
+Suppose Assumption 1 and 2 hold. The following inequalities on $s_{A}$, $s_{B}$ and $c$ hold:
+$\displaystyle 1\leq s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}},$ $\displaystyle\quad 1\leq s_{B}\leq\frac{M_{B}^{2}(1+\frac{1}{2}\ln(\kappa_{B}))}{1-\beta_{B}},$
+$\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\}.$
+where $M_{A}:=\max_{i\geq 0}\|A^{i}-A_{\infty}\|_{2}$, $M_{B}:=\max_{i\geq 0}\|B^{i}-B_{\infty}\|_{2}$.
 Proof.
@@
 3 Preliminary
-In this section, we provide a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms.
-Furthermore, we discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
-Push-Only and Pull-Only Algorithms Align with Centralized SGD.
-Most decentralized algorithms, such as decentralized SGD [36], decentralized gradient tracking [43, 14], (sub)gradient-Push [44], Push-Diging [7], (sub)gradient-Pull [9], and FROST [19], are designed to mimic centralized SGD.
-These algorithms can generally be formulated as:
+In this section, we present a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms.
+We also discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
+3.1 Push-Only and Pull-Only methods align with centralized SGD
+Most decentralized algorithms—such as decentralized SGD [36], gradient tracking [31, 32, 43, 44], gradient-Push [45], Push-DiGing [7], gradient-Pull [9], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting.
+These algorithms can generally be transformed into the following form:
 $\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-\alpha\bar{g}^{(k)}+\,\epsilon^{(k)}$ (5)
-where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}x_{i}^{(k)}$ is a convex combination of local decision variables, and $\bar{g}^{(k)}=\frac{1}{n}\sum_{i=1}^{n}\nabla F_{i}(x_{i}^{(k)},\xi_{i}^{(k)})$ represents the average gradient.
-Here, $\epsilon^{(k)}$ denotes an additional error term.
-For Push-Only and Pull-Only optimization algorithms, the extra error term is elaborately designed to be zero or to approach zero exponentially fast.
-Such design ensures that the trajectories of decision variables closely resemble those of centralized SGD.
-Given the sufficiently small extra error, the linear speedup result naturally follows within the same analytical framework used for centralized SGD.
-The following is a brief example of how to obtain the formulation (3).
-Example 1.
-Push-Diging [7]. The algorithm can be written as:
+where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}(x_{i}^{(k)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local optimization variables, with $v_{i}^{(k)}$ denoting the combination weights, and $\bar{g}^{(k)}=\frac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$ is the average stochastic gradient.
+The term $\epsilon^{(k)}$ captures additional error due to the decentralized nature of the algorithm.
+For Push-Only and Pull-Only optimization algorithms, the additional error term $\epsilon$ is carefully designed to be either zero or to decay exponentially fast.
+This design ensures that the trajectories of the local decision variables closely track those of centralized SGD.
+When the error is sufficiently small, the linear speedup result naturally follows from the same analytical framework used for centralized SGD.
+Below, we provide a brief example illustrating how to derive the formulation in (5).
+Example 1 (Push-Diging [7]).
+The Push-Diging algorithm iterates as follows:
 $\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A_{k}{\mathbf{x}}^{(k)}-\alpha V_{k+1}^{-1}{\mathbf{y}}^{(k)},$ (6a)
 $\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (6b)
-where $V_{k}=\mathrm{Diag}(v_{k})=\mathrm{Diag}(B^{k}\mathds{1})$ is a diagonal matrix and $A_{k}=V_{k+1}^{-1}BV_{k}$.
-By left-multiplying $n^{-1}v_{k+1}^{\top}$ on both sides of (4a) and omitting some calculation details, we will obtain
+where $v_{k}=B^{k}\mathds{1}$, $V_{k}=\mathrm{Diag}(v_{k})$ is a diagonal matrix, $A_{k}=V_{k+1}^{-1}BV_{k}$, and ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient defined in the notation section.
+By left-multiplying both sides of (6a) by $n^{-1}v_{k+1}^{\top}$, we obtain
 $n^{-1}v_{k+1}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}.$
-By taking $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k+1)}$, this is exactly the form of (3).
-Figure 1: Distributed logistic regression on synthetic dataset.
-We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005.
+Letting $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}$, this update exactly matches the form in (5).
+Example 2 (FROST [18, 19]).
+FROST/PullDiag-GT algorithm only relies on a row-stochastic matrix $A$. Its recursion can be written as follows:
+$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (7a)
+$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=A{\mathbf{y}}^{(k)}+D_{k+1}^{-1}{\mathbf{g}}^{(k+1)}-D_{k}^{-1}{\mathbf{g}}^{(k)},$ (7b)
+where $D_{k}=\mathrm{Diag}(A^{k})$ is a diagonal matrix, ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$.
+By left-multiplying both sides of (7a) by $n^{-1}\pi_{A}^{\top}$, we obtain
+$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}+\epsilon^{(k)},$
+where $\epsilon^{(k)}=\alpha n^{-1}\left(\mathds{1}^{\top}-\pi_{A}^{\top}D_{k}^{-1}\right){\mathbf{g}}^{(k)}$ converges to zero exponentially fast [Lemma 12, [39]].
+Letting $\hat{x}^{(k)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, this update matches the form  (5).
+Figure 1: The extra error does not vanish.
 The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant.
-For detailed settings, see 36, Appendix C.
-Push-Pull Analysis Suffers from Extra Error Terms.
-Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update.
-The recursion of Push-Pull is
+The experiment is a logistic regression on synthetic dataset.
+We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005, and detailed settings is referred to Appendix C.
+3.2 Push-Pull suffers from extra error terms
+Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
 $\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (8a)
 $\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (8b)
-[…]
-$\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$
-where $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$ is the projected decision variable, $c=n\pi_{A}^{\top}\pi_{B}$ is a constant.
+where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the topology, $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$, ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$.
+Left-multiplying both sides of (8b) by $\mathds{1}^{\top}$, we achieve
+$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(k+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(k)}+\mathds{1}^{\top}{\mathbf{g}}^{(k+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k+1)},$ (9)
+where the last equality holds due to the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Next, we left-multiply both sides of (8a) by $\pi_{A}^{\top}$, which yields
+$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}$
+$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}.$
+Let $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, and recall from (9) that $\mathds{1}^{\top}{\mathbf{y}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. Then, we have
+$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$ (10)
+where $c=n\pi_{A}^{\top}\pi_{B}$ is a constant and $\bar{g}^{(k)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$.
 We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
-This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\bar{g}^{(k)}$ is the dominating factor.
-But here, the error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|$ introduces additional noise that may dominate the noise in $\bar{g}^{(k)}$.
-For certain specific topologies, the noise in the error term can either be ignored or absorbed.
-For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull and demonstrates the linear speedup result on this particular topology.
-In their setting, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, which leads to the expression $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$.
-By setting the loss function at the first node to zero, they obtain the relationship $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$, which simplifies the process of achieving linear speedup.
-Unfortunately, for general topologies, the error term lacks exploitable structure, making direct analysis overly conservative.
-This leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result.
-Consequently, the central challenge becomes: How can we accurately characterize the noise within this error term?
+This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\|\bar{g}^{(k)}\|^{2}$ is the dominating factor.
+But here, both the squared error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|^{2}$ and the cross product $\left\langle\bar{g}^{(k)},\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\right\rangle$ can introduce additional noise that may dominate the noise in $\bar{g}^{(k)}$.
+For certain specific topologies, the error term can be ignored or absorbed.
+For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull (BTPP) and demonstrates the linear speedup result on this particular topology.
+In BTPP, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, leading to $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$.
+By setting the loss function at the first node to zero, $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$ is achieved, which simplifies the process of achieving linear speedup.
+Another case is when $A$ is doubly-stochastic; we have $\pi_{A}^{\top}(I-B_{\infty})=0$, which removes the error term.
+Unfortunately, for general topologies, the error term lacks exploitable structure.
+This usually leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result.
+Therefore, we recognize the central challenge in the analysis of stochastic Push-Pull: How to estimate the noise level in error terms more accurately?
 We tackle this key question in the following section.
@@
 Using the gradient tracking recursion (8b), we discover the following relationship about extra error terms: for any integer $k\geq 0,m\geq 1$:
-$\displaystyle\sum_{i=1}^{m}\Delta_{y}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}).$ (6)
-Equation (6) reveals a contraction-type recursion governing the sum of consecutive error terms.
-Informally, if we restrict our analysis to terms involving $\sigma^{2}$, we observe that $\|{\mathbf{g}}^{(k+m)}-{\mathbf{g}}^{(k)}\|F^{2}={\mathcal{O}}(\sigma^{2})$.
-This contraction property further implies that the aggregated term $\|\sum{i=1}^{m}\Delta_{y}^{(k+i)}\|_{F}^{2}$ maintains the same ${\mathcal{O}}(\sigma^{2})$ scaling.
-Consequently, we can deduce that individual terms satisfy $\|\Delta_{y}^{(k)}\|_{F}^{2}={\mathcal{O}}(\frac{\sigma^{2}}{m})$.
-This scaling relationship becomes particularly significant when $m$ is sufficiently large: in this regime, the $\sigma^{2}$-dependent component of the extra error term becomes negligible.
-As a result, our analysis preserves the crucial linear speedup property.
-To formalize this logic, we start from the following contraction relationship.
+$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (11)
+To understand this equation, we propose the following important supporting lemma:
 Lemma 3.
-For any integers $k\geq 0$ and $m\geq 1$, we have
-$\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (7)
-Furthermore, we can define a ”noiseless” version for the right-hand side, i.e.
-$H_{m,k}:=\pi_{A}^{\top}\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right){\mathbf{y}}^{(k)}+\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),$
-and the distance therein can be bounded by:
-$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\varepsilon^{(k+i)}-H_{m,k}\|_{F}^{2}]\leq 2ns_{B}^{2}\sigma^{2}.$ (8)
-The proof of Lemma 3 is provided in Appendix B.1.
-This lemma implies that the summation of the $m$ error terms can be represented as $H_{m,k}$ plus a bounded noise term.
-Moreover, the term $H_{m,k}$ exclusively involves difference of deterministic gradients and a single instance of ${\mathbf{y}}^{(k)}$, making it straightforward to estimate its magnitude.
-To effectively leverage the result of Lemma 3, we introduce an $m$-step descent lemma.
-This new lemma follows by an estimation on $H_{m,k}$ and applying the standard $L$-smoothness descent lemma to the following recursion:
+For any $l\geq 0$ and any matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$.
+For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
+$\displaystyle\mathbb{E}_{k}[\|H\|_{F}^{2}]\leq n\sigma^{2}\sum_{i=0}^{l}\|M^{(i)}\|_{2}^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{l}M^{(i)}\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}].$
+Proof.
+This is a natural corollary of Assumption 3. ∎
+Now we turn to the noise estimation for $\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$. A common analysis using Cauchy’s inequality typically provides
+$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim m^{2}\sigma^{2}.$ (12)
+Here, we use $\lesssim$ to absorb matrix-related constants and focus on $\sigma^{2}$-related terms.
+Nevertheless, analysis using Equation (11) and Lemma 3 can be significantly better.
+Proposition 2 ensures that $\textstyle\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+\|\sum_{j=1}^{m-1}B^{m-j-1}-B_{\infty}\|_{2}^{2}$ is finite.
+Therefore, applying Lemma 3 to Equation (11) tells us
+$\displaystyle\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\sigma^{2}$ (13)
+$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
+$\displaystyle\lesssim\sigma^{2}+\alpha^{2}L^{2}m^{2}\sigma^{2}.$
+Estimate from (13) sufficiently exploits the structure of gradient tracking, thus exhibiting ${\mathcal{O}}(m^{2})$ improvement compared to a naive estimate (12).
+To effectively leverage the result of  (13), we introduce an $m$-step descent lemma.
+It comes from applying the standard descent lemma to the following recursion:
 $\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}.$ (14)
 Lemma 4 (m-step Descent Lemma).
-Suppose Assumption 34 and Proposition 1 hold. For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$. The following inequality holds for any integers $k\geq 0,m\geq 1$:
-$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (10)
+Suppose Assumption 34 and Proposition 1 hold.
+For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$.
+When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for any integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$:
+$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (15)
 $\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}$
-$\displaystyle\quad+\frac{L}{16m}\Delta_{1}^{(k)}+\frac{L}{16m}\Delta_{2}^{(k)}+{3\alpha^{2}Ls_{B}^{2}}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}],$
+$\displaystyle\quad+\frac{c\alpha L^{2}}{2}\Delta_{1}^{(k)}+\frac{c\alpha L^{2}}{2}\Delta_{2}^{(k)}+\frac{3\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
 where
-$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
-Remarkably, if we divide both sides of (10) by $\frac{c\alpha m}{4}$, the first noise term simplifies to $\frac{8\alpha cL}{n}\sigma^{2}$, which corresponds to the standard linear speedup term.
-The second noise term simplifies to $\frac{8\alpha ns_{B}^{2}}{cm}\sigma^{2}$.
-When we select some $m$ larger than $c^{-2}n^{2}s_{B}^{2}$, the second noise term should have a smaller order than the first term, thereby ensuring the achievement of the standard linear speedup, which verifies our logic before.
-Next, we try to estimate $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$.
-Lemma 5 (Consensus Error).
-Suppose Assumption 34 and Proposition 1 hold. With $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{3c}{16mL},\frac{1}{6s_{B}L}\}$, we have
-$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}.$
-where $\Delta_{x}:={\mathbf{x}}^{(k)}-A_{\infty}{\mathbf{x}}^{(k)}$, $\Delta_{y}^{(k)}:={\mathbf{y}}^{(k)}-B_{\infty}{\mathbf{y}}^{(k)}$.
-Lemma 5 transfers the complicated error terms to commonly-seen consensus error terms. Now, we can apply techniques similar to [38] to bound $\Delta_{x}$ and $\Delta_{y}$.
-Lemma 6.
-Suppose Assumption 34 and Proposition 1 hold. For any $T\geq 1$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
-$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2},$
-and
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}].$
-Now we are ready to state the main result of this paper.
+$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
+The remaining errors are estimated in the next lemma.
+Lemma 5 (Remaining Errors).
+Suppose Assumption 34 and Proposition 1 hold.
+With $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\}$ and $m\geq C_{m}$, we have
+$\displaystyle\quad\frac{c\alpha L^{2}}{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq\frac{\alpha c}{4}\sum_{t=0}^{mK-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{\alpha cm}{4}(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
+where we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$, and $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants.
+$s_{B^{m}}$ is the measure defined for $B^{m}$, see (3).
+Generally, we have $s_{B^{m}}\leq s_{B}$, and $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}$ for large $m$.
+We are ready to state the main result of this paper.
 Theorem 1.
-Suppose Assumption 34 and Proposition 1 hold.
-For given positive integers $m$ and $K$, we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$ and $T:=mK$.
-Let $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$.
-When the learning rate $\alpha$ satisfies $\alpha\leq\min\{\frac{1}{8cL},\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L}\}$, we have
-$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$
-where $C_{1}$ is a constant determined by $A$ and $B$. Furthermore, if we choose
-$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\},$
-we will obtain that
-$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{C_{\Delta}L\Delta}{T},$ (11)
-where
-$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
-is a constant decided by $A$ and $B$.
+Suppose Assumption 34 and Proposition 1 hold. For given positive integers $m$ and $K$, we define $T:=mK$. When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq C_{m}$, we have
+$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (16)
+where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. Furthermore, if we choose
+$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
+$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
+we reach that
+$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]$ $\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$ (17)
+$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right).$
+Proof.
+See Appendix B.3
 Comparison with existing analyses.
-Theorem 1 provides the first finite-time non-convex convergence result for Push-Pull, while previous analyses focus either on the strongly convex case [45, 40] or on variants of Push-Pull relying on specific topologies [5].
-The first term $\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary strongly connected topologies.
-Impact of $m$.
-In our analysis, we consider a $m$-step gradient descent.
-The number $m$ effectively controls the $\sigma^{2}$-related term in the extra error.
-A common question is can we choose an arbitrarily large $m$ in this analysis.
-We can do that, but choosing a larger $m$ will force us to select a smaller learning rate, which finally leads to a larger transient time.
-To sum up, number $m$ provides a chance for tradeoff between the noise level and the transient time.
-Transient time.
-Our theorem provides a coarse estimate for the transient time, which is of ${\mathcal{O}}(nC_{\Delta}^{2})$.
-We do not recommend using this to estimate the real transient time, but it still reflects the impact of topology.
-For example, if we use a directed ring graph where $s_{A}=s_{B}={\mathcal{O}}(n)$, $c={\mathcal{O}}(1)$, the transient time can be ${\mathcal{O}}(n^{17})$.
-If we use a directed exponential graph [3] where $s_{A}=s_{B}={\mathcal{O}}(1)$, $c={\mathcal{O}}(1)$, the transient time will be ${\mathcal{O}}(n^{5})$.
+Theorem 1 provides the first finite-time non-convex convergence result for stochastic Push-Pull, while previous analyses focus either on the strongly convex case [46, 40] or on variants of Push-Pull relying on specific topologies [5].
+The first term $\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary topologies satisfying Proposition 2.
+Impact of $m$: In our analysis, we consider $m$-step gradient descent.
+The parameter $m$ effectively controls the terms related to $\sigma^{2}$ from additional errors, as illustrated by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (16).
+$m$ cannot be arbitrarily large because larger values of $m$ restrict the range of the learning rate, resulting in slower convergence and longer transient times in the analysis.
+In summary, the parameter $m$ allows for a tradeoff between noise level and transient time.
+Transient time. Our theorem is able to provide a coarse estimate for the transient time, which is
+$\max\{n^{2}c^{-4}s_{B}^{8},n^{7}c^{-4}s_{A}^{2}s_{B}^{3},nc^{-4}s_{B}^{4}\}.$
+The order of $n$ is overestimated because we frequently use $\|\pi_{A}\|\leq 1$ and $\|\pi_{B}\|\leq 1$, leading to a loss of $n$ when $A$ or $B$ is doubly-stochastic.
 5 Experiment
@@
 where $\Delta^{(i)}\in\mathbb{R}^{n\times d}$ are arbitrary matrices, $s_{A}$ is defined as in (2). Furthermore, if Assumption 1 and 2 hold, we have
-$\displaystyle s_{A}:=\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$
-where $\beta_{A}$, $\kappa_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
+$\displaystyle s_{A}:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
+where $\beta_{A}$, $\kappa_{A},M_{A}$ are defined in Proposition 2. The row-stochastic matrix $A$ and the subscript $A$ can be replaced by a column-stochastic matrix $B$ without changing the lemma.
 Proof.
@@
 Second, we want to prove that
-$\displaystyle s_{A}\leq\frac{1+\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}.$ (14)
+$\displaystyle s_{A}\leq\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$ (20)
 Towards this end, we define $M_{A}:=\max\{1,\max_{k\geq 0}\|A^{k}-A_{\infty}\|_{2}\}$.
@@
 If $p>0$, let $x=\min\{\lfloor p\rfloor,k\}+1-p\in[0,1)$, (20) is simplified to
-$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}}).$
-By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$. This finishes the proof of (14).
+$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}(x+p+\frac{\beta_{A}^{x}}{1-\beta_{A}})\leq M_{A}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
+By the definition of $p$, we have $p\leq\frac{\frac{1}{2}\ln(\kappa_{A})}{1-\beta_{A}}$.
+This completes the estimation of $\sum_{i=0}^{\infty}\|A^{k-i}-A_{\infty}\|_{2}$.
+For the summation of the squared norm, we can use the same $p$ and obtain the upper bound without modifying the previous steps:
+$\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}^{2}\leq M_{A}^{2}(p+\frac{1}{1-\beta_{A}})=\frac{M_{A}^{2}(1+\frac{1}{2}\ln(\kappa_{A}))}{1-\beta_{A}}.$
+This finishes the proof of (20).
 Finally, to obtain (18), we use Jensen’s inequality. For any positive numbers $a_{0},a_{1},\ldots,a_{k}$ satisfying $\sum_{i=0}^{k}a_{i}=1$, we have
@@
 B.1 Proof of Lemma 4
-For $1\displaystyle 1\leq i\leq m-$ 1, consider
-$\displaystyle(I-B_{\infty}){\mathbf{y}}^{(k+i)}$ $\displaystyle=(I-B_{\infty})(B{\mathbf{y}}^{(k+i-1)}+{\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k+i-1)})$
-$\displaystyle=(B-B_{\infty}){\mathbf{y}}^{(k+i-1)}+(I-B_{\infty}){\mathbf{g}}^{(k+i)}-(I-B_{\infty}){\mathbf{g}}^{(k+i-1)}$
-Sum it up for $1\displaystyle 1\leq i\leq m-$ 1, we obtain that
-$(I-B_{\infty})\sum_{i=1}^{m-1}{\mathbf{y}}^{(k+i)}=(B-B_{\infty})\sum_{i=0}^{m-2}{\mathbf{y}}^{(k+i)}+(I-B_{\infty}){\mathbf{g}}^{(k+m-1)}-(I-B_{\infty}){\mathbf{g}}^{(k)}.$
-Similarly, consider $(B-B_{\infty}){\mathbf{y}}^{(k+i)}$, $2\displaystyle 1\leq i\leq m-$ 2, we have
-$(B-B_{\infty})\sum_{i=1}^{m-2}{\mathbf{y}}^{(k+i)}=(B^{2}-B_{\infty})\sum_{i=0}^{m-3}{\mathbf{y}}^{(k+i)}+(B-B_{\infty}){\mathbf{g}}^{(k+m-2)}-(B-B_{\infty}){\mathbf{g}}^{(k)}.$
-Repeat this procedure for $m-3,m-4,\ldots,1$ and sum them up, we finally arrive at
-$\sum_{i=0}^{m-1}(I-B_{\infty}){\mathbf{y}}^{(k+i)}=\sum_{i=0}^{m-1}(B^{i}-B_{\infty}){\mathbf{y}}^{(k)}+\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
-which is just the first statement. Next, note that the difference between $\sum_{i=0}^{m-1}\varepsilon^{(k+i)}$ and $H_{m,k}$ equals to:
-$\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))-\pi_{A}^{\top}\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})).$
-By taking the squared norm, we obtain that:
-$\displaystyle\quad\sum_{i=1}^{m-1}\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}$
-$\displaystyle+\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}+\text{Inner Product Terms}.$
-When we take the expectation conditioned on ${\mathcal{F}}_{k}$, all the inner product terms become zero. For example, for any $1\displaystyle 0\leq i<j\leq m-$ 1,
-$\displaystyle\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\right\rangle\right]$
-$\displaystyle=\mathbb{E}_{k}[\mathbb{E}_{k+j}[\text{This inner product}]]$
-$\displaystyle=\mathbb{E}_{k}\left[\left\langle(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})),0\right\rangle\right]=\mathbb{E}[0]=0.$
-Therefore, we only need to estimate the squared terms. By the Cauchy-Schwarz inequality, and $\|\pi_{A}\|\leq 1$, we have
-$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}^{\top}(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|^{2}]$
-$\displaystyle\leq\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(B^{m-i-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)}))\|_{F}^{2}]$
-$\displaystyle\leq\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k+i)}-\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}]$
-$\displaystyle\leq n\sigma^{2}(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}\leq ns_{B}^{2}\sigma^{2},$
-and
-$\displaystyle\quad\mathbb{E}_{k}[\|\pi_{A}^{\top}(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$
-$\displaystyle\leq\mathbb{E}_{k}[\|\pi_{A}\|^{2}\|(\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-$\displaystyle\leq\|\sum_{i=1}^{m-1}(B^{m-i-1}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|{\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
-$\displaystyle\leq(\sum_{i=1}^{m-1}\|B^{m-i-1}-B_{\infty}\|_{2})^{2}n\sigma^{2}\leq ns_{B}^{2}\sigma^{2}.$
-By combining the two parts, we finish the proof. $\square$
-B.2 Proof of Lemma 4
 For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x^{(k)}_{i})$. According to Assumption 3, we have
-$\mathbb{E}[\bar{g}^{(k+i)}]=\mathbb{E}[\mathbb{E}_{k}[\bar{g}^{(k+i)}]]=\mathbb{E}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
+$\mathbb{E}_{k}[\bar{g}^{(k+i)}]=\mathbb{E}_{k}[\mathbb{E}_{k+i}[\bar{g}^{(k+i)}]]=\mathbb{E}_{k}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
 Note that
@@
 $\displaystyle\quad+c^{2}\alpha^{2}L\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}+\alpha^{2}L\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}.$
-By taking expectation (from time 00 0), we obtain that
-$\displaystyle\mathbb{E}[f(\hat{x}^{(k+m)})-f(\hat{x}^{(k)})]\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (18)
-$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
-$\displaystyle\quad+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
+By taking expectation conditioned on ${\mathcal{F}}_{k}$, we obtain that
+$\displaystyle\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]-f(\hat{x}^{(k)})\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}_{k}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (24)
+$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]$
+$\displaystyle\quad+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
 The first inner product term in (24) can be bounded by:
@@
 where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness. The first quartic term in (24) can be bounded by:
-$\mathbb{E}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
-where we use the fact that $\bar{g}^{(k+i)}$ are conditional unbiased estimate and
-$\mathbb{E}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
+$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}]\leq n^{-1}m\sigma^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}],$
+where we use the fact that $\bar{g}^{(k+i)}$ are conditionally unbiased estimates and
+$\mathbb{E}_{k}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
 Plug these estimations into (24), we obtain that
-$\displaystyle\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (19)
-$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
-$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
+$\displaystyle\frac{\alpha cm}{2}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (25)
+$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]+\frac{\alpha^{2}c^{2}mL}{n}\sigma^{2}+c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$
+$\displaystyle-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]+\alpha^{2}L\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}].$
 Next, we try to simplify the last quadratic term in (25) . Notice that
@@
 the quadratic term can be bounded as:
-$\displaystyle\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (20)
-$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
+$\displaystyle\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$ (26)
+$\displaystyle\leq 2\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+2\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}].$
 The first term here can be estimated using the fact that $\|\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\|_{2}^{2}\leq s_{B}^{2}$. For the second term in (26), according to Assumption 3, we have
-$\displaystyle\quad\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (21)
-$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
-$\displaystyle\quad+\mathbb{E}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+$\displaystyle\quad\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$ (27)
+$\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
+$\displaystyle\quad+\mathbb{E}_{k}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
 $\displaystyle\leq n\sigma^{2}\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+n\sigma^{2}\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}^{2}$
-$\displaystyle\quad+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-The inner product can produce a similar term. According to Lemma 3,
-$\mathbb{E}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}[H_{m}({\mathbf{y}}^{(k)},\nabla f({\mathbf{x}}^{(k)},\ldots,\nabla f({\mathbf{x}}^{(k+m-1)}))].$
-Therefore, we have
-$\displaystyle\quad-\alpha\mathbb{E}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (22)
-$\displaystyle=-\alpha\mathbb{E}\left[\left\langle\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})),\nabla f(\hat{x}^{(k)})\right\rangle\right]$
-$\displaystyle\quad+\mathbb{E}\left[\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right){\mathbf{y}}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$
-$\displaystyle\leq\frac{1}{cmL}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-$\displaystyle\quad+\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\frac{\alpha cm}{2}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}].$
-Plug (20) and (22) into (19), we obtain that
-$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (23)
+$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+$\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+The inner product can produce a similar term:
+$\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}_{k}[H_{m,k}].$
+Using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
+$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]=-\alpha\left\langle\mathbb{E}_{k}[H_{m,k}],\nabla f(\hat{x}^{(k)})\right\rangle$ (28)
+$\displaystyle=-\alpha\left\langle\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))],\nabla f(\hat{x}^{(k)})\right\rangle$
+$\displaystyle\quad+\alpha\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle$
+[…]
+$\displaystyle\quad+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
+$\displaystyle\leq\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]+\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}$
+$\displaystyle\quad+\frac{2\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
+where the last inequality uses $\|\pi_{A}\|\leq 1$ and $\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\|_{2}\leq s_{B}$. Plug (26) and (28) into (25), we obtain that
+$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]-c^{2}\alpha^{2}L\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$ (29)
+$\displaystyle\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
+$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+(2\alpha^{2}Ls_{B}^{2}+\frac{2\alpha s_{B}^{2}}{cm})\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle+(2\alpha^{2}L+\frac{2\alpha}{cm})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+For $c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
+$\displaystyle\quad\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}].$ (30)
+Take (30) into (29), and suppose $\alpha\leq\frac{1}{4cmL}$, we obtain that
+$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (31)
 $\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-For $2c^{2}\alpha^{2}L\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]$, we have the following estimate:
-$\displaystyle\quad\mathbb{E}[\|\sum_{i=0}^{m-1}\overline{\nabla f}^{(k+i)}\|^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}].$ (24)
-Take (24) into (23), we obtain that
-$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (25)
+$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+We can further apply Cauchy’s inequality on the last term in (31), i.e.
+$\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
+$\displaystyle\leq\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
+$\displaystyle\leq s_{B}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
+Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (31) can be further simplified to
+$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
 $\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-We can further apply Jensen inequality on the last term in (25), i.e.
-$\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
-$\displaystyle\leq(\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2})^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}$
-$\displaystyle\leq s_{B}^{2}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
-Therefore, (25) can be further simplified to
-$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c(1-4cm\alpha L)}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
-$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{c\alpha L^{2}}{2}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle+(2\alpha^{2}L+\frac{1}{cmL})s_{B}^{2}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
-By taking $\alpha\leq\min\{\frac{1}{8cmL},\frac{c}{8n^{2}s_{B}^{2}L}\}$, this can be further simplified to
-$\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (26)
-$\displaystyle\leq\mathbb{E}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]+\frac{L}{16m}\sum_{i=0}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+3\alpha^{2}Ls_{B}^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{L}{16m}\sum_{j=1}^{m-1}\mathbb{E}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
+$\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{j=1}^{k}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
 which finishes the proof.
-B.3 Proof of Lemma 5
-We consider how to measure the difference of ${\mathbf{x}}$. Use the fact that
-$\displaystyle{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
-$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)}),$
-We decompose ${\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}$ and apply the Cauchy-Schwarz inequality:
-$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]\leq 3\left(\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\right)\|\Delta_{x}^{(k)}\|_{F}^{2}$ (27)
-$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
-$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})\|_{F}^{2}]$
-$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$
-$\displaystyle\quad+3\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-$\displaystyle\leq 3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}+3\alpha^{2}s_{B}^{2}L^{2}\Delta_{2}^{(k)},$
-where the last inequality comes from Cauchy-Schwarz inequality and $L$-smoothness.
-For ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, similarly, we decompose it into
-$\displaystyle{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$ $\displaystyle=(A^{i}-A_{\infty})\Delta_{x}^{(k)}-\alpha\left(\sum_{j=0}^{i-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}$
-$\displaystyle\quad-\alpha\sum_{j=1}^{i-1}(B^{i-j-1}-B_{\infty})({\mathbf{g}}^{(k+i)}-{\mathbf{g}}^{(k)})+\Delta_{x}^{(k)},$
-which finally produces
-$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]\leq 4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (28)
-$\displaystyle\quad+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4\alpha^{2}s_{B}^{2}L^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]+4m\|\Delta_{x}^{(k)}\|_{F}^{2}.$
-Inequalities (27) and (28) tell us
-$\displaystyle\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}$
-where $v_{1}=3s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+3\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+6\alpha^{2}mns_{B}^{2}\sigma^{2}$, $v_{2}=4s_{A}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}ms_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+8\alpha^{2}mns_{B}^{2}\sigma^{2}+4m\|\Delta_{x}^{(k)}\|_{F}^{2}$ are positive numbers.
-To proceed on, we assume $\alpha\leq\frac{1}{6s_{B}L}$, so that the inversion of the $2\times 2$ coefficient matrix satisfies
-$\displaystyle\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\leq{\begin{pmatrix}1&-3\alpha^{2}s_{B}^{2}L^{2}\\ 0&1-4\alpha^{2}s_{B}^{2}L^{2}\end{pmatrix}}^{-1}\leq\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}.$
-Therefore, we have
-$\displaystyle{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}\mathbb{E}_{k}[\Delta_{1}^{(k)}]\\ \mathbb{E}_{k}[\Delta_{2}^{(k)}]\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{16m}\\ \frac{L}{16m}\end{pmatrix}}^{\top}\begin{pmatrix}1&0.1\\ 0&1.5\end{pmatrix}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}\leq{\begin{pmatrix}\frac{L}{10m}\\ \frac{L}{10m}\end{pmatrix}}^{\top}\begin{pmatrix}v_{1}\\ v_{2}\end{pmatrix}.$
-This gives that
-$\displaystyle\quad\frac{L}{16m}\mathbb{E}_{k}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]\leq\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2},$
-which finishes the proof of Lemma 5.
-B.4 Proof of Lemma 6
-In this section, we provide some inequalities for the consensus error $\Delta_{x}$ and $\Delta_{y}$. These inequalities will only be used in the final part of proof of our main theorem.
-Lemma 8 (Consensus of x).
-Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, we have
-$\displaystyle\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k+1)}\|^{2}]$ $\displaystyle\leq 2\alpha^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{y}^{(k)}\|^{2}]$
-$\displaystyle\quad+2\alpha^{2}n^{2}s_{A}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|^{2}]+2n\alpha^{2}s_{A}^{2}(T+1)\sigma^{2}$
+B.2 Proof of Lemma 5
+In this section, we consider how to measure the difference of ${\mathbf{x}}$. To be more specific, we mainly consider the relationships between the following 44 44 terms:
+$\Delta_{1}^{(k)},\Delta_{2}^{(k)},\mathbb{E}_{k}[\|\Delta_{x}^{(k)}\|_{F}^{2}],\Delta_{y}^{(k)},\quad k\in[m^{*}].$
+Lemma 7 will give upper bounds of $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ using other terms, while Lemma
+Lemma 7.
+Suppose Proposition 2 and Assumption 34 hold.
+When $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold for any positive integers $k$ and $m$.
+$\displaystyle\Delta_{1}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (32)
+$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
+$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (33)
+$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
 Proof.
-First note that
-$\displaystyle\Delta_{x}^{(k+1)}$ $\displaystyle=(I-A_{\infty}){\mathbf{x}}^{(k+1)}=(A-A_{\infty}){\mathbf{x}}^{(k)}-(I-A_{\infty}){\mathbf{y}}^{(k)}$
-$\displaystyle=\ldots=\sum_{i=0}^{k}(A^{k-i}-A_{\infty}){\mathbf{y}}^{(i)}$
-$\displaystyle=\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta_{y}^{(i)}+\sum_{i=0}^{k}(A^{k-i}-A_{\infty})n\pi_{B}\bar{g}^{(i)}$
+Define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1}$, when $j\geq 1$. We have
+$\displaystyle\quad{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}-(A^{i}-A_{\infty})\Delta_{x}^{(k)}=-\alpha\sum_{j=0}^{i-1}A^{i-j-1}\Delta_{y}^{(k+j)}-\alpha\sum_{j=0}^{i-1}A^{i-j-1}B_{\infty}{\mathbf{y}}^{(k+j)}$
+$\displaystyle=-\alpha\sum_{j=1}^{i-1}A^{i-j-1}(B^{j}-B_{\infty})\Delta_{y}^{(k)}-\alpha\sum_{l=1}^{i-1}\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}({\mathbf{g}}^{(k+l)}-{\mathbf{g}}^{(k)})$
+$\displaystyle\quad\quad-\alpha\sum_{j=0}^{i-1}(A^{i-j-1}-A_{\infty})n\pi_{B}\bar{g}^{(k+j)}-\alpha c\sum_{j=0}^{i-1}\bar{g}^{(k+j)}.$
+Therefore, we can apply Cauchy’s inequality and sum it up from $i=1$ to $m-1$:
+$\displaystyle\quad\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq 5\sum_{i=1}^{m-1}\|A^{i}-A_{\infty}\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+5\alpha^{2}\sum_{i=1}^{m-1}\|\sum_{j=1}^{i-1}A^{i-j-1}(B^{j}-B_{\infty})\|_{2}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}$
+$\displaystyle\quad+5\alpha^{2}\sum_{i=1}^{m-1}(\sum_{l=1}^{i-1}\|\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}\|_{2}^{2}+\|\sum_{l=1}^{i-1}\sum_{j=0}^{i-l-1}A^{i-j-l-1}\delta_{B}^{(j)}\|_{2}^{2})n\sigma^{2}$
+$\displaystyle\quad+5\alpha^{2}\sum_{i=1}^{m-1}\mathbb{E}_{k}[\|\sum_{j=l}^{i-1}A^{i-j-1}\delta_{B}^{(j-l)}(\nabla f({\mathbf{x}}^{(k+l)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+$\displaystyle\quad+5\alpha^{2}n^{2}s_{A}^{2}\sum_{i=1}^{m-1}\|\overline{\nabla f}^{(k+j)}\|+5\alpha^{2}n\sum_{i=1}^{m-1}\sum_{j=0}^{i-1}\|A^{i-j-1}-A_{\infty}\|_{2}^{2}$
+$\displaystyle\quad+5c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\|\overline{\nabla f}^{(k+j)}\|^{2}+5c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$
+$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+ms_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
+$\displaystyle\quad+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
+where in the first inequality, the $\sigma^{2}$ terms are separated using Lemma 3, and the second inequality comes from rolling sum.
+$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}$
+$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}$
+For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$. This gives
+$\displaystyle\quad\Delta_{2}^{(k)}\leq 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
+$\displaystyle\quad+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
+Easy to prove that $\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\leq 2s_{A}+2mn$. When $\alpha\leq\frac{1}{5N_{8}L}$, we have
+$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (34)
+$\displaystyle\quad+10\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+10\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
+Plug (34) into the estimate for $\Delta_{1}^{(k)}$, it is not difficult to verify that $\Delta_{1}^{(k)}$ could use the same upper bound as (34).
+Next, we provide some inequalities for the consensus error $\Delta_{x}^{(k)}$ and $\Delta_{y}^{(k)}$, $k\in[m^{*}]$.
+For $k\in[m^{*}]$, we have
+$\displaystyle\Delta_{y}^{(k+m)}=(B^{m}-B_{\infty})\Delta_{y}^{(k)}+\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$ (35)
+Define
+$\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)}).$
+Note that it can be rewritten as
+$\displaystyle\Delta_{3}^{(k)}$ $\displaystyle=(I-B_{\infty}){\mathbf{g}}^{(k+m)}+(B-I){\mathbf{g}}^{(k+m-1}+(B^{2}-B){\mathbf{g}}^{(k+m-2)}$
+$\displaystyle\quad+\ldots+(B^{m-1}-B^{m-2}){\mathbf{g}}^{(k+1)}-(B^{m-1}-B_{\infty}){\mathbf{g}}^{(k)},$
+according to Lemma 3, we have
+$\displaystyle\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$
+where
+$\displaystyle h_{m,k}$ $\displaystyle:=\sum_{j=1}^{m}\delta_{B}^{(m-j)}\nabla f({\mathbf{x}}^{(k+j)})-(B^{m-1}-B_{\infty})\nabla f({\mathbf{x}}^{(k)}),$
+$\displaystyle q_{B}$ $\displaystyle=\sum_{j=1}^{m}\|\delta_{B}^{(m-j)}\|_{2}^{2}+\|B^{m-1}-B_{\infty}\|_{2}^{2}\leq 2s_{B}.$
+For $h_{m,k}$, we can rearrange it as
+$\displaystyle h_{m,k}$ $\displaystyle=\sum_{j=1}^{m}\delta_{B}^{(m-j)}(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})).$
+By applying the rolling sum lemma, we have
+$\displaystyle\mathbb{E}_{k}[\|h_{m,k}\|_{F}^{2}]$ $\displaystyle\leq\sum_{t=0}^{m}\|\delta_{B}^{(t)}\|_{2}^{2}\sum_{j=1}^{m}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
+$\displaystyle\leq 2s_{B}L^{2}\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=2s_{B}L^{2}\Delta_{2}^{(k)}.$
+Combining the estimates above, we reach that
+$\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq 2ns_{B}\sigma^{2}+2s_{B}L^{2}\Delta_{2}^{(k)}.$
+To proceed on, we iterate (35) and obtain that
+$\displaystyle\Delta_{y}^{(k+m)}=(B^{m+k}-B_{\infty}){\mathbf{g}}^{(0)}+\sum_{j=0}^{k/m}(B^{k-jm}-B_{\infty})\Delta_{3}^{(jm)}.$
 Using the rolling sum lemma, we have
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{x}^{(k+1)}\|_{F}^{2}\leq 2s_{A}^{2}\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}+2n^{2}s_{A}^{2}\|\pi_{B}\|^{2}\sum_{k=0}^{T}\|\bar{g}^{(k)}\|_{F}^{2}.$
-Note that $\mathbb{E}[\|\bar{g}^{(k)}\|_{F}^{2}]\leq n^{-1}\sigma^{2}+\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$ and $\|\pi_{B}\|\leq 1$, we finish the proof. ∎
-Lemma 9 (Consensus of y).
-Suppose Assumption 43 and Proposition 1 hold. For any $T\geq 0$, when $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, we have
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
-Proof.
-$\displaystyle\Delta_{y}^{(k+1)}$ $\displaystyle=(B-B_{\infty})\Delta_{y}^{(k)}+(I-B_{\infty})({\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)})$
-$\displaystyle=\ldots=(B^{k+1}-B_{\infty}){\mathbf{g}}^{(0)}+\sum_{i=0}^{k}(B^{k-i}-B_{\infty})({\mathbf{g}}^{(i+1)}-{\mathbf{g}}^{(i)})$
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k+m)}\|_{F}^{2}]\leq 2\sum_{k\in[m^{*}]}\mathbb{E}[\|(B^{k}-B_{\infty}){\mathbf{g}}^{(0)}\|_{F}^{2}]+2s_{B^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]]$
+$\displaystyle\quad\leq s_{B^{m}}(n\sigma^{2}+2nL\Delta)+4ns_{B}s_{B^{m}}^{2}K\sigma^{2}+4s_{B}s_{B^{m}}^{2}L^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}].$
+Note that
+$\mathbb{E}[\|\Delta_{y}^{(0)}\|_{F}^{2}]=\mathbb{E}[\|(I-B_{\infty}){\mathbf{g}}^{(0)}\|_{F}^{2}]\leq 2(n\sigma^{2}+2nL\Delta),$
+we combine coefficients and reach that
+$\displaystyle\sum_{k\in[m^{*}]}\|\Delta_{y}^{(k)}\|_{F}^{2}$ $\displaystyle\leq 7s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+6s_{B^{m}}nL\Delta+4s_{B}s_{B^{m}}^{2}L^{2}\sum_{k\in[m^{*}]}\Delta_{2}^{(k)}.$
+Plug in (33), we have
+$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}$
+$\displaystyle\,\,+6s_{B^{m}}nL\Delta+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
+When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}s_{B}s_{B^{m}}$, we have $1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2}\geq 0.5$.
+and the coefficient of $Kn\sigma^{2}$ becomes smaller than $8s_{B}s_{B^{m}}^{2}$.
+Thus,
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq 16s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+12s_{B^{m}}nL\Delta$ (36)
+$\displaystyle\quad+320s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+160c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
+By taking this back to (33), we have
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]$ $\displaystyle\leq 40mn(1+320\alpha^{2}L^{2}ms_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]$ (37)
+$\displaystyle\quad+480\alpha^{2}ms_{B^{m}}nL\Delta+(640ms_{B}^{2}s_{B^{m}}^{2}+20c^{2}m^{2}n^{-2})\alpha^{2}Kn\sigma^{2}$
+$\displaystyle\quad+20c^{2}\alpha^{2}m^{2}(1+160\alpha^{2}L^{2}ms_{B}^{2}s_{B^{m}}^{2})\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
+$\displaystyle\leq 50mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+480\alpha^{2}ms_{B^{m}}nL\Delta$
+$\displaystyle\quad+21c^{2}m^{2}n^{-1}\alpha^{2}K\sigma^{2}+21c^{2}\alpha^{2}m^{2}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
+where the second inequality uses $\alpha\leq\frac{1}{10cmL}$ and $m\geq 4c^{-2}s_{B}^{2}s_{B^{m}}^{2}n^{2}$.
+For $k\in[m^{*}]$, we have
+$\displaystyle\Delta_{x}^{(k+m)}=(A^{m}-A_{\infty})\Delta_{x}^{(k)}-\alpha\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}.$ (38)
+Define
+$\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)},$
+then we can iterate (38) and obtain that
+$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(k)}.$
+To further analyze $\Delta_{4}^{(k)}$, we expand it as
+$\displaystyle\Delta_{4}^{(k)}$ $\displaystyle=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\Delta_{y}^{(k+i)}+n\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\pi_{B}\bar{g}^{(k+i)}$
+$\displaystyle=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})(B^{i}-B_{\infty})\Delta_{y}^{(k)}+n\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\pi_{B}\bar{g}^{(k+i)}$
+$\displaystyle\quad+\sum_{j=i}^{m-1}(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$
+Note that
+$\displaystyle\|\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})(B^{i}-B_{\infty})\|_{2}^{2}$ $\displaystyle\leq s_{A}s_{B},\,\,\sum_{i=0}^{m-1}\|(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}\leq 2s_{A}s_{B},$
+$\displaystyle\sum_{i=0}^{m-1}\|(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}^{2}$ $\displaystyle\leq 2s_{A}s_{B},\,\,\|\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\delta_{B}^{(j-i)}\|_{2}^{2}\leq 2s_{A}s_{B}.$
+Therefore, according to Lemma 3, we can separate noiseless and noisy terms in $\sum_{i=0}^{m-1}\pi_{B}\bar{g}^{(k+i)}$ and $\sum_{j=i}^{m-1}(A^{m-i}-A_{\infty})(B^{j-i}-B^{j-i+1})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)})$, then apply Jensen’s inequality.
+This gives
+$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
+$\displaystyle\quad+6s_{A}s_{B}\sum_{j=i}^{m-1}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]+12s_{A}s_{B}n\sigma^{2}$
+$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]$
+$\displaystyle\quad+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
 Using the rolling sum lemma, we have
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq s_{B}^{2}\|{\mathbf{g}}^{(0)}\|+s_{B}^{2}\sum_{k=0}^{T}\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}$
-By Cauchy-Schwarz Inequality and our assumption for bounded noise, we have
-$\displaystyle\mathbb{E}[\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 6n\sigma^{2}+3\mathbb{E}[\|\nabla f({\mathbf{x}}^{(k+1)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
-$\displaystyle\leq 6n\sigma^{2}+3L^{2}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]$
-Besides, $\mathbb{E}[\|{\mathbf{g}}^{(0)}\|_{F}^{2}]\leq n\sigma^{2}+\|\nabla f({\mathbf{x}}^{(0)})\|_{F}^{2}\leq n\sigma^{2}+2nL\Delta$. Therefore, we have
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7)\sigma^{2}+2ns_{B}^{2}L\Delta+3s_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$ (29)
-Since
-$\displaystyle\quad\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=\mathbb{E}[\|(A-I){\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)}\|_{F}^{2}]$ (30)
-$\displaystyle=\mathbb{E}[\|(A-I)\Delta_{x}^{(k)}-\alpha\Delta_{y}^{(k)}+n\alpha\pi_{B}\bar{g}^{(k)}\|_{F}^{2}]$
-$\displaystyle\leq 3\|A-I\|_{2}^{2}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+3\alpha^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+3n\alpha^{2}\sigma^{2}+3n^{2}\alpha^{2}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}],$
-and $\|A-I\|_{2}^{2}\leq 2n$, we can apply (30) to substitute the last term in (29). This produces
-$\displaystyle(1-9\alpha^{2}L^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T)\sigma^{2}$
-$\displaystyle\quad+9ns_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+9\alpha^{2}L^{2}n^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]+2ns_{B}^{2}L\Delta$
-Using the conclusion of Lemma 8, this can be further transformed to
-$\displaystyle\quad(1-9\alpha^{2}L^{2}s_{B}^{2}-18\alpha^{2}L^{2}ns_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$
-$\displaystyle\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T+18\alpha^{2}L^{2}ns_{A}^{2}T)\sigma^{2}+2ns_{B}^{2}L\Delta$
-$\displaystyle\quad+(9\alpha^{2}L^{2}n^{2}s_{B}^{2}+18\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
-When $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, this can be simplified to
-$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
-B.5 Proof of Theorem 1
-Plug Lemma 5 into Lemma 4, we obtain that
-$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (31)
-$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\left(\frac{2\alpha^{2}c^{2}mL}{n}+4\alpha^{2}Lns_{B}^{2}\right)\sigma^{2}$
-$\displaystyle\quad+\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}.$
-When $\alpha\leq\frac{1}{8cmL}$, $\frac{\alpha c(1-4\alpha mcL)}{2}$ can be replaced with $\frac{c\alpha}{4}$.
-Sum up (31) for $k\in[m^{*}]$, take full expectation and divide both sides by $\frac{c\alpha T}{4}$ we obtain that
-$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (32)
-$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{4\alpha Lns_{B}^{2}}{cm}\right)\sigma^{2}$
-$\displaystyle\quad+\frac{4s_{A}^{2}L}{cm\alpha T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+\frac{16\alpha Ls_{B}^{2}}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
-Since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{x}^{(k+1)}\|_{F}^{2}$, we can plug in Lemma 8 into (32), and obtain that
-$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (33)
-$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}\right)\sigma^{2}$
-$\displaystyle\quad+\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{16\alpha L(s_{A}^{4}+s_{B}^{2})}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
-Similarly, since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$, we further plug in Lemma 9 into (33), and obtain that
-$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (34)
-$\displaystyle\leq\frac{4\Delta}{c\alpha T}+C_{\sigma}\sigma^{2}+\frac{128ns_{B}^{2}\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}+C_{f}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
+$\displaystyle\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]]$
+$\displaystyle\leq 15\alpha^{2}ns_{A}s_{B}s_{A^{m}}^{2}K\sigma^{2}+3\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
+$\displaystyle\quad+6\alpha^{2}L^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]+3\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}\|\Delta_{y}^{(k)}\|_{F}^{2}$
+Note that we start from consensual $x$, so $\Delta_{x}^{(0)}=0$. Plug in (37) and (36) we have
+$\displaystyle\quad(1-480\alpha^{2}mL^{2}s_{A}s_{B}s_{A^{m}}^{2}n-960\alpha^{2}L^{2}ms_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2}n)\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq(15+126c^{2}\alpha^{2}L^{2}m^{2}n^{-2}+48s_{B}s_{B^{m}}^{2})\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}nK\sigma^{2}$
+$\displaystyle\quad+(3+(126+480s_{B}s_{B^{m}}^{2})c^{2}\alpha^{2}L^{2}m^{2})\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
+$\displaystyle\quad+(36+2880\alpha^{2}L^{2})s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}nL\Delta$
+Using $\alpha\leq\min\{\frac{1}{10\sqrt{s_{B}s_{B^{m}}^{2}}cmL},\frac{1}{6L}\}$ and $m\geq 20c^{-2}ns_{A}s_{B}s_{A^{m}}^{2}$, the coefficients can be simplified to
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq 64\alpha^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2}nK\sigma^{2}$ (39)
+$\displaystyle\quad+7\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+416\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}nL\Delta$
+Taking (39) back to (36), we obtain that
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(1+1280\alpha^{2}L^{2}mn^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2})16s_{B}s_{B^{m}}^{2}Kn\sigma^{2}$
+$\displaystyle\quad+(12+133120\alpha^{2}L^{2}s_{A}s_{B}^{2}s_{B^{m}}s_{A^{m}}^{2})s_{B^{m}}nL\Delta$
+$\displaystyle\quad+160(c^{2}m^{2}s_{B}s_{B^{m}}^{2}+140\alpha^{2}L^{2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}s_{B^{m}}^{2})\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
+When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}$, we have
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 32s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+26s_{B^{m}}nL\Delta$ (40)
+$\displaystyle\quad+170c^{2}m^{2}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
+Plug (39) into (37), we obtain that
+$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]\leq(21c^{2}m^{2}n^{-2}+3200mns_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2})\alpha^{2}nK\sigma^{2}$ (41)
+$\displaystyle\quad+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+(21c^{2}m+350n^{3}s_{A}s_{A^{m}}^{2})\alpha^{2}m\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
+$\displaystyle\leq 181c^{2}m^{2}n^{-1}\alpha^{2}K\sigma^{2}+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+38\alpha^{2}c^{2}m^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}],$
+where the second inequality uses $m\geq 20c^{-2}n^{3}s_{A}s_{A^{m}}^{2}s_{B}^{2}s_{B^{m}}^{2}$.
+Now we are ready to estimate the summation of error terms and prove Lemma 5.
+Remember that $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ share the same upper bound, we utilize (41), (40) and obtain that
+$\displaystyle\frac{c\alpha L^{2}}{2}\quad\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq\frac{\alpha cm}{4}C_{\sigma,1}K\sigma^{2}+\frac{c\alpha}{4}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{c\alpha L\Delta}{4m},$
 where
-$\displaystyle C_{\sigma}$ $\displaystyle=\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}+\frac{480\alpha Lns_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cm},$
-$\displaystyle C_{f}$ $\displaystyle=\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}+\frac{2000\alpha^{3}L^{3}n^{3}s_{A}^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cmT}.$
-When $m\geq 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$, we have $C_{\sigma}\leq\frac{30c\alpha L}{n}$ and $C_{f}\leq\frac{2c\alpha L}{5T}+\frac{100c\alpha^{3}L^{3}n^{3}s_{A}^{2}}{T}$.
-When we further set $\alpha\leq\min\{\frac{1}{8cL},\frac{1}{4ns_{A}L}\}$, we have $C_{f}\leq\frac{1}{T}$.
-Under these conditions, (34) can be simplified to
-$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$ (35)
-Finally, to satisfy all the constraints on $m$ and $\alpha$, we can set $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$, and
-$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\}.$
-Plugging this into (35) gives that
-$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+C_{\Delta}\frac{L\Delta}{T},$
-where
-$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
-is a constant decided by $A$ and $B$.
+$\displaystyle C_{\sigma,1}$ $\displaystyle=724c^{2}mn^{-1}\alpha^{2}L^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$
+$\displaystyle C_{f,1}$ $\displaystyle=152c^{2}\alpha^{2}L^{2}m^{2}+2040m\alpha^{2}L^{2}s_{B}^{3}s_{B^{m}}^{2}$
+$\displaystyle C_{\Delta,1}$ $\displaystyle=85120s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}L^{2}m^{2}n^{2}+\frac{312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
+When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq c^{-2}30s_{B}^{1.5}$, we have
+$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
+This finishes the proof.
+B.3 Proof of Theorem 1
+Sum up Lemma 4 for $k\in[m^{*}]$, then plug in Lemma 5, we obtain that
+$\displaystyle\quad\frac{\alpha cm}{4}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}(1-C_{f,1})\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (42)
+$\displaystyle\leq f(x^{(0)})-f(\hat{x}^{T})+\frac{\alpha cm}{4}\left(\frac{8\alpha cL}{n}+\frac{8\alpha Lns_{B}^{2}}{cm}+C_{\sigma,1}\right)K\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
+Divide both sides by $\frac{\alpha cmK}{4}$ and utilize $C_{f,1}\leq 1$, $f(x^{(0)})-f(\hat{x}^{T})\leq\Delta$, $m\geq c^{-2}n^{2}s_{B}^{2}$ and $C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$, $T=mK$, we have
+$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (43)
+Inequality (43) holds when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}},\frac{1}{6L}\}$ and $m$ is larger than some matrix-related constants.
+To obtain a standard linear speedup result, we set
+$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
+$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
+giving that
+$\displaystyle\quad\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]$
+$\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$
+$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)$
 Appendix C Experiment Details

### v3

@@
 The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts.
-In this paper, we study the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm over directed graphs (digraphs).
-Unlike methods that rely solely on row-stochastic or column-stochastic mixing matrices, Push-Pull avoids nonlinear correction and has shown superior empirical performance across a variety of settings.
-However, its theoretical analysis remains challenging, and the linear speedup property has not been generally established—revealing a significant gap between empirical success and limited theoretical understanding.
-To bridge this gap, we propose a novel analysis framework and prove that Push-Pull achieves linear speedup over arbitrary strongly connected digraphs.
-Our results provide the comprehensive theoretical understanding for stochastic Push-Pull, aligning its theory with empirical performance.
+This paper studies the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm for directed graphs (digraphs).
+Unlike methods that rely on Push-Only or Pull-Only communications, Push-Pull avoids nonlinear correction and has shown superior empirical performance across various settings.
+Despite its success, whether Push-Pull achieves linear speedup has remained an open question, revealing a significant gap between its empirical performance and theoretical understanding.
+To bridge this gap, we propose a novel multi-step descent analysis framework and prove that Push-Pull indeed achieves linear speedup over arbitrary strongly connected digraphs.
+Our results provide a comprehensive theoretical understanding of the stochastic Push-Pull method, finally aligning its theoretical convergence with empirical performance.
 Code: https://github.com/pkumelon/PushPull.
 1 Introduction
-Scaling machine learning tasks to large datasets and models necessitates efficient distributed computation across multiple nodes.
-In distributed optimization, a network of $n$ nodes cooperates to solve
-$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F(x;\xi_{i})].$ (1)
-Here, $\xi_{i}$ denotes a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function, which may be nonconvex.
-During optimization, each node $i$ can access its own local gradient $\nabla F_{i}(x;\xi_{i})$, but must communicate to obtain gradient information from the other nodes.
-Data heterogeneity is typically present in (1), meaning the local distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ differ across nodes.
-Consequently, the loss functions $f_{i}(x)$ and $f_{j}(x)$ are generally not equal, i.e., $f_{i}(x)\neq f_{j}(x)$.
-Parallel stochastic gradient descent (SGD) is a widely adopted approach in distributed optimization, where local gradients are globally averaged using either a central server [1] or Ring All-Reduce communication primitives [2].
-However, such global synchronization incurs substantial bandwidth overhead and high latency as the number of nodes increases, thereby hurting the scalability of the optimization process.
-To address these limitations, decentralized SGD has recently been proposed to solve problem (1) without requiring global gradient aggregation.
-In this setting, each node independently maintains and updates its model by communicating only with immediate neighbors, significantly reducing communication overhead.
-Beyond efficiency, decentralized algorithms offer improved robustness to node and link failures, maintaining convergence as long as the underlying communication graph remains connected.
-In contrast, centralized SGD breaks down if the central server crashes.
-In this paper, we consider a general decentralized setting where the network topology is modeled as a strongly connected directed graph (digraph), which naturally encompasses undirected networks as a special case.
-Compared to undirected graphs, digraphs are particularly well-suited for ad hoc decentralized systems, such as heterogeneous robotic networks that operate under asynchronous, one-sided communication and random node activation.
-More recently, digraphs have shown great potential in distributed deep learning, where certain classes of directed networks exhibit both topological sparsity and strong algebraic connectivity.
-These properties make them especially attractive for the distributed training of large-scale models.
-Notable examples include the exponential graph [3], EquiTopo [4], and B-ary tree [5].
+Scaling modern machine learning tasks to massive datasets and increasingly large models requires efficient distributed computation across multiple nodes.
+In distributed optimization, a network of $n$ nodes cooperatively solves
+$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}(x;\xi_{i})].$ (1)
+where each local function $f_{i}$ is defined by a data distribution $\mathcal{D}_{i}$.
+Each node has access only to stochastic gradients $\nabla F_{i}(x;\xi_{i})$ from its local data, while communication is required to aggregate information across the network.
+In practice, the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are heterogeneous, so the local losses $f_{i}$ generally differ.
+This heterogeneity poses significant challenges to both the design and analysis of distributed algorithms.
+A natural baseline for solving (1) is centralized stochastic gradient descent (SGD), where gradients are aggregated globally either through a central server [1] or collective communication primitives such as Ring All-Reduce [2].
+While centralized SGD enjoys strong theoretical guarantees—including the celebrated linear speedup property in iteration complexity—its reliance on global averaging leads to significant communication overhead and creates a single point of failure at the server.
+These drawbacks limit scalability in large, heterogeneous systems.
+Decentralized SGD has emerged as a compelling alternative [3, 4].
+In this setting, each node maintains its own local model and communicates only with its neighbors.
+This localized communication substantially reduces bandwidth usage and alleviates latency bottlenecks.
+Moreover, decentralized methods are inherently robust to failures: as long as the communication graph remains connected, the network can continue to train without dependence on a single master node.
+In this paper, we consider decentralized optimization over directed graphs (digraphs).
+Directed communication naturally models many real-world systems, including heterogeneous robotic swarms with asymmetric links, sensor networks with one-sided message passing, and distributed deep learning environments where communication is constrained by bandwidth asymmetry.
+Directed networks are also attractive because certain classes of sparse digraphs, such as exponential graphs [5], EquiTopo [6], and B-ary trees [7], achieve both scalability and strong connectivity.
 1.1 Decentralized algorithms over digraphs
-In any connected network, the topology can be characterized by a mixing matrix.
+In a connected network, the topology can be characterized by a mixing matrix.
 This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance.
@@
 When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [8, 13] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity.
-More advanced methods such as EXTRA-Push [13], DEXTRA [14], ADD-OPT [15], and Push-DIGing [7] improve convergence speed by mitigating the effects of data heterogeneity.
+More advanced methods such as EXTRA-Push [14], DEXTRA [15], ADD-OPT [16], and Push-DIGing [9] improve convergence rate by mitigating the effects of data heterogeneity.
 Stochastic variants, including SGD-Push [5] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
-The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting.
-Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods.
-Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting.
-Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
+The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only methods.
+Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [11] serves as the foundation for Pull-Only methods.
+Reference [11] adapted the distributed gradient descent algorithm to the Pull-Only setting.
+Subsequently, gradient tracking techniques were extended to this framework by [18, 19], while momentum-based Pull-Only gradient tracking methods were developed in [20, 21].
 The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23].
@@
 As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices.
-[25] extended the Push-Pull method to stochastic optimization settings, while [5] improved its convergence rate on B-ary tree network topologies.
+Reference [25, 26] extended the Push-Pull method to stochastic optimization settings, while [7] improved its convergence rate on B-ary tree network topologies.
 Accelerated or constrained variant of the Push-Pull method have also been explored in [27, 28].
 1.2 Linear speedup in iteration complexity
-Linear speedup refers to the property of a distributed algorithm whereby the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes.
-For example, centralized SGD achieves an $\epsilon$-accurate solution in $\mathcal{O}\left(\textstyle\frac{1}{n\epsilon^{2}}\right)$ iterations , where $n$ denotes the number of nodes.
-This scaling behavior demonstrates that centralized SGD achieves linear speedup with respect to the number of nodes.
-The linear speedup property is crucial for establishing the advantage of distributed algorithms over their single-node counterparts.
-Without this property, the practical utility of a distributed algorithm may be limited, as it would offer little benefit from using multiple computing nodes.
-The linear speedup property of decentralized algorithms over undirected graphs is well established.
-The first such analysis for decentralized SGD was presented in [28], demonstrating that decentralized SGD can achieve the same asymptotic convergence rate as its centralized counterpart.
-Building on this foundation, a variety of advanced decentralized algorithms—such as Exact Diffusion/NIDS [29, 30] and Gradient Tracking methods [31, 32, 33]—are proven to achieve linear speedup [34, 29, 35].
-In addition, [36] provides a linear speedup analysis for decentralized SGD under time-varying undirected network topologies.
-Recent literature has increasingly focused on establishing the linear speedup property of decentralized algorithms over digraphs.
-For Push-Only algorithms that employ column-stochastic mixing matrices, linear speedup has been formally proven in [16, 37].
-Moreover, a recent study [38] proposes methods that achieve the optimal iteration complexity in this setting.
-For Pull-Only algorithms that employ row-stochastic mixing matrices, convergence has been extensively studied under various settings [9, 18, 19, 20, 21]; however, these analyses do not establish the linear speedup property.
-The main challenge lies in the fact that using a row-stochastic matrix alone introduces a deviation in the descent direction from the globally averaged gradient.
-With a novel analysis framework addressing this issue, the linear speedup property of Pull-Only algorithms was established only recently in [39].
-These results reveal the effectiveness of Push-Only and Pull-Only algorithms over digraphs.
-However, although Push-Pull algorithms typically achieve significantly better empirical performance than Push-Only and Pull-Only methods, their linear speedup property has not been generally established.
-Existing theoretical analyses of Push-Pull remain limited, primarily focusing on strongly convex loss functions in deterministic settings [23, 40, 22, 5].
-In the nonconvex and stochastic setting, neither linear speedup nor even basic convergence guarantees have been formally established for Push-Pull.
-The most relevant theoretical result to date is presented in [5], which introduces a variant called B-ary Tree Push-Pull (BTPP).
-While BTPP is shown to achieve linear speedup, the result relies on the specialized B-ary tree topology, and the analysis does not generalize to general network topologies.
+The linear speedup is a central property for distributed optimization.
+It states that the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes $n$.
+Centralized SGD achieves this scaling, converging in $\mathcal{O}(1/(n\epsilon^{2}))$ iterations.
+Without this property, the benefit of parallelization is limited, as increasing the number of nodes would not yield proportional gains in iteration complexity.
+For decentralized algorithms on undirected networks, linear speedup is well established.
+The first rigorous analysis for decentralized SGD was given in [29].
+Building on this work, a range of advanced algorithms—including Exact Diffusion/NIDS [30, 31] and gradient tracking methods [32, 33, 34, 9]—have been proven to achieve linear speedup [35].
+Time-varying network extensions are also addressed in [36].
+For digraphs, theoretical understandings are less explored.
+For Push-Only algorithms with column-stochastic mixing, linear speedup has been established in [5, 37], with subsequent refinements achieving optimal iteration complexity in [38].
+For Pull-Only algorithms with row-stochastic mixing, convergence has been extensively analyzed under various settings [11, 18, 19, 20, 21], but these studies did not establish linear speedup because row-stochastic matrices inherently introduce a deviation from the globally averaged gradient.
+Only recently, linear speedup for Pull-Only algorithms was proven in [39].
+These advances reveal the effectiveness of both Push-Only and Pull-Only methods over digraphs.
+Despite this progress, establishing the linear speedup property for Push-Pull algorithms—those combining both row- and column-stochastic matrices—has remained an open problem.
+While Push-Pull methods consistently outperform Push-Only and Pull-Only variants in practice, their theoretical understanding remains limited.
+Existing analyses focus mostly on strongly convex settings [23, 25, 22, 7], and provide no guarantees for nonconvex or stochastic optimization.
+The most relevant result to date is [7], which demonstrates linear speedup for Push-Pull under a specialized B-ary tree topology, but the analysis cannot be extended to general network structures.
+A recent work [40] establishes guarantees for Push-Pull in a specific non-convex federated learning setting, but does not prove the linear speedup property.
+Table 1 provides a summary of existing results on decentralized algorithms over digraphs.
 1.3 Motivation and main results
-Building on the above discussion, the existing theoretical convergence analyses for Push-Pull methods remain rather limited, underscoring a clear discrepancy between the weak theoretical guarantees and their strong empirical performance.
-This gap gives rise to the following two open questions:
+The above discussion highlights a striking gap between theory and practice: Push-Pull algorithms consistently outperform Push-Only and Pull-Only methods in empirical studies, yet their scalability theory is far less developed.
+This gap motivates two fundamental questions:
   • Q1.
@@
   • Q2.
-    Can Push-Pull achieve linear speedup over a broader class of network topologies—for instance, all strongly connected digraphs?
-Addressing these two questions is crucial for establishing a theoretical foundation that supports the effectiveness of Push-Pull methods.
-Regarding Question Q1, we demonstrate that both Push-Only and Pull-Only algorithms closely align with centralized SGD; specifically, after a careful numerical transformation, their variable updates closely resemble those of centralized SGD.
-This structural similarity explains why the linear speedup property can be readily established for Push-Only and Pull-Only methods.
-In contrast, Push-Pull is not designed to align with centralized SGD and consistently exhibits a non-vanishing deviation from the centralized trajectory.
-Consequently, the linear speedup analysis for Push-Pull is substantially more challenging and fundamentally distinct from that of Push-Only and Pull-Only algorithms.
-For Question Q2, we establish a linear speedup convergence result for the Push-Pull method.
-To the best of our knowledge, this is the first such result for Push-Pull over all strongly connected digraphs under standard assumptions, whereas existing analyses of stochastic Push-Pull [5, 41] are restricted to tree-structured topologies.
-By introducing a novel analysis framework that addresses the misalignment between Push-Pull and centralized SGD, this paper bridges the gap between the convergence theory and the strong empirical performance of the Push-Pull method.
+    Can push–pull achieve linear speedup on general strongly connected digraphs under standard assumptions?
+In this work, we address both questions.
+Regarding Question Q1, we show that Push-Only and Pull-Only can be reformulated to align closely with centralized SGD, making their linear speedup analysis relatively straightforward.
+In contrast, Push-Pull exhibits a persistent, non-vanishing deviation from centralized SGD dynamics, rendering the conventional single-step descent analysis framework ineffective.
+For Question Q2, we introduce a novel multi-step descent analysis framework that aggregates errors over blocks of iterations.
+This framework uncovers a telescoping structure in the gradient-tracking recursion: intermediate gradient noise cancels out, while boundary terms are exponentially attenuated by mixing dynamics.
+Consequently, the residual error remains bounded at a constant scale, enabling us to establish the desired $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity for Push-Pull.
+To the best of our knowledge, this paper presents the first general proof that Push-Pull achieves linear speedup on arbitrary strongly connected digraphs in the stochastic nonconvex setting.
+In particular, our multi-step descent analysis framework not only bridges the gap between theory and practice for Push-Pull but also introduces techniques that may be of independent interest for the broader study of decentralized algorithms.
+A summary of existing results is provided in Table 1.
+Table 1: Convergence rate comparison of representative decentralized algorithms in non-convex and stochastic settings over directed graphs.
+“Conv.
+Rate” refers to the asymptotic rate as $T\to\infty$; “Iter.
+Complex.” denotes the number of iteration requried to achieve $\epsilon$-accuracy; “S.C.” denotes “Strongly Connected”; “R-Sto.”, “C-Sto.” and “R/C Alt.” refer to “Row-Stochastic”, “Column-Stochastic” and “Row/Column Alternate”, respectively.
+Note that References [25, 26] study the Push-Pull algorithm under strongly convex objectives, rather than the non-convex setting.
+$\sigma$ denotes gradient noise.
+Algorithms
+Conv.
+Rate
+Iter.
+Complex.
+Topology
+Mixing
+Matrix
+Centralized SGD $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ Star graph -
+Pull-Diag-GT [39] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R-Sto.
+Push-Diging [38] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph C-Sto.
+Push-Pull [25] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
+Push-Pull [26] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
+Push-Pull [40] $\mathcal{O}(\frac{\sigma}{\sqrt{T}})$ $\mathcal{O}(\frac{\sigma^{2}}{\epsilon^{2}})$ S.C. digraph R/C Alt.
+Push-Pull [7] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ B-ary Tree R/C Alt.
+Push-Pull (Ours) $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R/C Alt.
 The remainder of the paper is organized as follows.
-Section 2 introduces necessary notations, assumptions, and measures.
-Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods, and offers insights into the derivation of linear speedup convergence.
-In Section 4, we present the key components of our analysis and state the main convergence theorem.
-Section 5 provides extensive numerical experiments that demonstrate the ability of the Push-Pull method to achieve linear speedup.
+Section 2 introduces the notations, assumptions, and performance metrics used throughout the paper.
+Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods and provides insights into the derivation of their linear speedup properties.
+Section 4 presents our novel multi-step descent analysis framework and explains the intuition behind its role in enabling linear speedup analysis.
+Section 5 formally develops the supporting lemmas and states the main convergence theorem.
+Section 6 reports extensive numerical experiments that validate the ability of the Push-Pull method to achieve linear speedup.
 Finally, Section 7 concludes the paper.
-2 Notations, Assumptions and Measures
-This section introduces notations, assumptions, and measures that will be used throughout the paper.
+2 Notations, Algorithms, Assumptions and Metrics
+This section introduces notations, assumptions, and metrics that will be used throughout the paper.
 Notations.
-The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ is the set of directed edges.
+The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ the set of directed edges.
 An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$.
-A directed graph is said to be strongly connected if there exists a directed path from any node to every other node in the graph.
-We say that $\mathcal{G}=(\mathcal{N},\mathcal{E})$ is the underlying topology of the weight matrix $A=[a_{ij}]$ if $A$ is entrywise nonnegative and $a_{ij}>0$ if and only if $(j,i)\in\mathcal{E}$.
-Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector.
-Set $[n]$ denotes $\{1,2,\ldots,n\}$.
-For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ as its diagonal entries.
-We use matrix $A$ to denote the row-stochastic matrix (i.e., $A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix (i.e., $B^{\top}\mathds{1}=\mathds{1}$).
-The Perron vectors [42] of matrices $A$ and $B$, if they exist, are denoted by $\pi_{A}$ and $\pi_{B}$, respectively.
-We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$.
+The graph is strongly connected if a directed path exists between every pair of nodes.
+We say $\mathcal{G}$ is the underlying topology of a nonnegative weight matrix $A=[a_{ij}]$ if $a_{ij}>0$ precisely when $(j,i)\in\mathcal{E}$.
+Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector and $[n]:=\{1,2,\ldots,n\}$.
+For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ on its diagonal, while $\min v$ and $\max v$ denote its smallest and largest entries, respectively.
+We let $A$ denote a row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and $B$ a column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$), with Perron vectors [41] $\pi_{A}$ and $\pi_{B}$, respectively.
+Define $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$.
 Additionally, we define $n\times d$ matrices
-$\displaystyle{\mathbf{x}}^{(k)}$ $\displaystyle:=[({x}_{1}^{(k)})^{\top};({x}_{2}^{(k)})^{\top};\cdots;({x}_{n}^{(k)})^{\top}]$
-$\displaystyle{\mathbf{g}}^{(k)}$ $\displaystyle:=[\nabla F_{1}({x}_{1}^{(k)};\xi_{1}^{(k)})^{\top};\cdots;\nabla F_{n}({x}_{n}^{(k)};\xi_{n}^{(k)})^{\top}]$
-$\displaystyle\nabla f({\mathbf{x}}^{(k)})$ $\displaystyle:=[\nabla f_{1}({x}^{(k)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(k)}_{n})^{\top}]$
-[…]
+$\displaystyle{\mathbf{x}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[({x}_{1}^{(t)})^{\top}\hskip-1.42262pt;({x}_{2}^{(t)})^{\top}\hskip-1.42262pt;\cdots;({x}_{n}^{(t)})^{\top}\hskip-1.42262pt]$
+$\displaystyle\hskip-1.42262pt{\mathbf{g}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla F_{1}({x}_{1}^{(t)};\hskip-1.42262pt\xi_{1}^{(t)})^{\top}\hskip-1.42262pt;\hskip-1.42262pt\cdots\hskip-1.42262pt;\hskip-1.42262pt\nabla F_{n}({x}_{n}^{(t)};\xi_{n}^{(t)})^{\top}\hskip-1.42262pt]$
+$\displaystyle\nabla f({\mathbf{x}}^{(t)})$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla f_{1}({x}^{(t)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(t)}_{n})^{\top}$]
+where subscripts index nodes and superscripts index iterations.
+Let ${\mathcal{F}}_{t}$ be the filtration containing all information up to ${\mathbf{x}}^{(t)}$, excluding the stochastic gradients at iteration $t$.
+We use $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot\mid{\mathcal{F}}_{t}]$ for conditional expectation.
+Norms are denoted as follows: $\|\cdot\|$ for the vector $\ell_{2}$-norm, $\|\cdot\|_{F}$ for the Frobenius norm, and $\|\cdot\|_{2}:=\max_{\|v\|=1}\|Av\|$ for the matrix spectral norm.
+Push-Pull algorithm. The Push-Pull algorithm, initially proposed for deterministic decentralized optimization [22, 23], can be naturally extended to the stochastic gradient setting [25]:
+$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (2a)
+$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (2b)
+where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the underlying digraph.
+For initialization, we let $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$ and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$.
+The row-stochastic mixing matrix $A$ in (2a) facilitates consensus among the local variables maintained by each node, whereas the column-stochastic mixing matrix $B$ in (2b) enables the computation of a globally averaged gradient.
+In the absence of gradient noise, i.e., when ${\mathbf{g}}^{(t)}=\nabla f({\mathbf{x}}^{(t)})$, it has been shown in [22, 23] that ${\mathbf{x}}^{(t+1)}$ converges to ${\mathbf{x}}^{\star}=[(x^{\star})^{\top};\cdots;(x^{\star})^{\top}]\in\mathbb{R}^{n\times d}$ for strongly convex problems, where $x^{\star}$ is the solution to problem (1).
 Assumptions. We first specify the underlying graphs ($\mathcal{G}_{A}$, $\mathcal{G}_{B}$) and the corresponding weight matrices ($A$, $B$) used in the stochastic Push-Pull method.
 Assumption 1 (Communication Graph).
-Graphs ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ are strongly connected.
+Assume ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ to be strongly connected digraphs (see the definition in notations) with $n$ nodes.
 Assumption 2 (Weight Matrix).
@@
 $\pi_{A}^{\top}A=\pi_{A}^{\top},\quad B\pi_{B}=\pi_{B},\quad\|\pi_{A}\|_{1}=\|\pi_{B}\|_{1}=1,\quad\pi_{A}^{\top}\pi_{B}>0.$
-Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
-$\|A^{k}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{k},\quad\|B^{k}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{k},\quad\forall k\geq 0.$
+Besides, there exist constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
+$\|A^{t}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{t},\quad\|B^{t}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{t},\quad\forall k\geq 0.$ (3)
 Proof.
 The proof follows directly from the Perron-Frobenius theorem [41]. ∎
-Finally, we state standard assumptions regarding the stochastic gradients and function smoothness.
+We next state assumptions regarding stochastic gradients and function smoothness.
 Assumption 3 (Gradient Oracle).
 There exists a constant $\sigma\geq 0$ such that
-$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)})\quad\mbox{and}\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2}$
-for any $i\in[n]$ and $k\geq 0$.
+$\displaystyle\mathbb{E}_{t}[\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(t)})\quad\mbox{and}\quad\mathbb{E}_{t}[\|\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})-\nabla f_{i}(x_{i}^{(t)})\|_{F}^{2}]\leq\sigma^{2}$
+for any $i\in[n]$ and $t\geq 0$. In addition, we assume that stochastic gradients at different nodes are independent for any $t\geq 0$ and $1\leq i<j\leq n$.
 Assumption 4 (smoothness).
@@
 $\displaystyle\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|\quad\mbox{and}\quad f_{i}(x)-\inf_{x}f_{i}(x)\leq\Delta,\quad\forall i\in[n].$
-Measures.
-Various effective metrics have been proposed to capture the impact of digraphs on decentralized stochastic optimization.
-Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$.
-Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{1/2}(B-B_{\infty})\Pi_{B}^{-1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$.
+Metrics.
+Various effective metrics have been proposed in literautre to capture the impact of digraphs on decentralized stochastic optimization.
+Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$ where $\pi_{A}$ and $\pi_{B}$ are defined in Proposition 1.
+Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{-1/2}(B-B_{\infty})\Pi_{B}^{1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$.
 Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
-However, this paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define the following constants:
-$\displaystyle s_{A}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\},$ (2)
-$\displaystyle s_{B}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\},$ (3)
+This paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define:
+$\displaystyle s_{A}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\big\},$ (4)
+$\displaystyle s_{B}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\big\},\vskip 5.69054pt$ (5)
 $\displaystyle c$ $\displaystyle:=n\,\pi_{A}^{\top}\pi_{B}.$ (6)
-Throughout this paper, we characterize the influence of the mixing matrices using only $s_{A}$, $s_{B}$, $c$, and $n$.
-A connection between these quantities and the previously discussed metrics is established in Proposition 2.
-Proposition 2 (Relation with other measures).
+By the exponential decay property (3), both $s_{A}$ and $s_{B}$ are finite.
+In this paper, the influence of the mixing matrices is described solely through $s_{A}$, $s_{B}$, $c$, and $n$.
+Their connection to previously introduced metrics is given in the next proposition.
+Proposition 2 (Relation with existing metrics).
 Suppose Assumption 1 and 2 hold. The following inequalities on $s_{A}$, $s_{B}$ and $c$ hold:
@@
 $\displaystyle n\max\{\min\pi_{A},\min\pi_{B}\}$ $\displaystyle\leq\,c\leq n\min\{\max\pi_{A},\max\pi_{B}\}.$
-where $M_{A}:=\max_{i\geq 0}\|A^{i}-A_{\infty}\|_{2}$, $M_{B}:=\max_{i\geq 0}\|B^{i}-B_{\infty}\|_{2}$.
+where $M_{A}:=\max_{t\geq 0}\|A^{t}-A_{\infty}\|_{2}$ and $M_{B}:=\max_{t\geq 0}\|B^{t}-B_{\infty}\|_{2}$ are finite numbers which can be proved to be no larger than $\sqrt{n}$.
 Proof.
 See Appendix A. ∎
-3 Preliminary
-In this section, we present a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms.
-We also discuss the limitations of existing analytical approaches and explain why they cannot be directly extended to the Push-Pull algorithm.
-3.1 Push-Only and Pull-Only methods align with centralized SGD
-Most decentralized algorithms—such as decentralized SGD [36], gradient tracking [31, 32, 43, 44], gradient-Push [45], Push-DiGing [7], gradient-Pull [9], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting.
-These algorithms can generally be transformed into the following form:
-$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-\alpha\bar{g}^{(k)}+\,\epsilon^{(k)}$ (5)
-where $\hat{x}^{(k)}=\sum_{i=1}^{n}v_{i}^{(k)}(x_{i}^{(k)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local optimization variables, with $v_{i}^{(k)}$ denoting the combination weights, and $\bar{g}^{(k)}=\frac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$ is the average stochastic gradient.
-The term $\epsilon^{(k)}$ captures additional error due to the decentralized nature of the algorithm.
-For Push-Only and Pull-Only optimization algorithms, the additional error term $\epsilon$ is carefully designed to be either zero or to decay exponentially fast.
-This design ensures that the trajectories of the local decision variables closely track those of centralized SGD.
-When the error is sufficiently small, the linear speedup result naturally follows from the same analytical framework used for centralized SGD.
-Below, we provide a brief example illustrating how to derive the formulation in (5).
+3 Intrinsic Differences between Push-Only, Pull-Only, and Push-Pull
+This section provides a concise overview of the linear speedup analysis for the Push-Only and Pull-Only algorithms, clarifying their key differences from Push-Pull.
+We also highlight the limitations of existing analyses and explain why they cannot establish the linear speedup property of Push-Pull.
+3.1 Push-Only and Pull-Only align with centralized SGD
+Many decentralized algorithms—such as decentralized SGD [36], gradient tracking [32, 33, 34], Push-Diging [9], Gradient-Pull [11], and FROST [19]—are designed to emulate centralized SGD in a decentralized setting.
+These algorithms can be reformulated as
+$\displaystyle\hat{x}^{(t+1)}=\underbrace{\hat{x}^{(t)}-\alpha\bar{g}^{(t)}}_{\text{centralized SGD}}+\epsilon^{(t)},$ (7)
+where $\hat{x}^{(t)}=\sum_{i=1}^{n}v_{i}^{(t)}(x_{i}^{(t)})^{\top}\in\mathbb{R}^{1\times d}$ is a convex combination of the local decision variables $\{x_{i}^{(t)}\}_{i=1}^{n}$, with $v_{i}^{(t)}$ denoting the combination weights.
+The term $\bar{g}^{(t)}=\tfrac{1}{n}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$ is the globally averaged stochastic gradient, and $\epsilon^{(t)}$ captures the residual error due to decentralization.
+When $\epsilon^{(t)}=0$, the recursion in (7) reduces to centralized SGD, whose linear speedup property is well established (see Table 1).
+For Push-Only and Pull-Only algorithms, the residual $\epsilon^{(t)}$ is either exactly zero or decays exponentially to zero.
+This ensures that the trajectory of the weighted average of local iterates closely tracks that of centralized SGD.
+Once the residual is sufficiently small, the linear speedup property follows directly from the standard analysis of centralized SGD.
+We illustrate this alignment with two examples.
 Example 1 (Push-DiGing [9]).
-The Push-Diging algorithm iterates as follows:
-$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A_{k}{\mathbf{x}}^{(k)}-\alpha V_{k+1}^{-1}{\mathbf{y}}^{(k)},$ (6a)
-$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (6b)
-where $v_{k}=B^{k}\mathds{1}$, $V_{k}=\mathrm{Diag}(v_{k})$ is a diagonal matrix, $A_{k}=V_{k+1}^{-1}BV_{k}$, and ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient defined in the notation section.
-By left-multiplying both sides of (6a) by $n^{-1}v_{k+1}^{\top}$, we obtain
-$n^{-1}v_{k+1}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}.$
-Letting $\hat{x}^{(k)}=n^{-1}v_{k}^{\top}{\mathbf{x}}^{(k)}$, this update exactly matches the form in (5).
+Push-DiGing is a Push-Only method, whose iterates are given by
+$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A_{t}{\mathbf{x}}^{(t)}-\alpha V_{t+1}^{-1}{\mathbf{y}}^{(t)},$ (8a)
+$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (8b)
+where $B$ is column-stochastic, $v_{t}=B^{t}\mathds{1}$, $V_{t}=\mathrm{Diag}(v_{t})$, $A_{t}=V_{t+1}^{-1}BV_{t}$, and ${\mathbf{g}}^{(t)}$ stacks local stochastic gradients.
+Left-multiplying (8a) by $n^{-1}v_{t+1}^{\top}$ yields
+$n^{-1}v_{t+1}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}.$
+Defining $\hat{x}^{(t)}=n^{-1}v_{t}^{\top}{\mathbf{x}}^{(t)}$, the recursion matches (7) with $\epsilon^{(t)}=0$.
 Example 2 (FROST [18, 19]).
-FROST/PullDiag-GT algorithm only relies on a row-stochastic matrix $A$. Its recursion can be written as follows:
-$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (7a)
-$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=A{\mathbf{y}}^{(k)}+D_{k+1}^{-1}{\mathbf{g}}^{(k+1)}-D_{k}^{-1}{\mathbf{g}}^{(k)},$ (7b)
-where $D_{k}=\mathrm{Diag}(A^{k})$ is a diagonal matrix, ${\mathbf{g}}^{(k)}$ is the stacked stochastic gradient and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$.
-By left-multiplying both sides of (7a) by $n^{-1}\pi_{A}^{\top}$, we obtain
-$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\bar{g}^{(k)}+\epsilon^{(k)},$
-where $\epsilon^{(k)}=\alpha n^{-1}\left(\mathds{1}^{\top}-\pi_{A}^{\top}D_{k}^{-1}\right){\mathbf{g}}^{(k)}$ converges to zero exponentially fast [Lemma 12, [39]].
-Letting $\hat{x}^{(k)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, this update matches the form  (5).
-Figure 1: The extra error does not vanish.
-The top three plots illustrate the convergence behavior across different network topologies, while the bottom three plots demonstrate that the size of extra error term rapidly stabilizes and persists as a significant constant.
-The experiment is a logistic regression on synthetic dataset.
-We use stochastic gradients and a constant learning rate 0.0050.005 0.0050.005, and detailed settings is referred to Appendix C.
-3.2 Push-Pull suffers from extra error terms
-Unlike the aforementioned algorithms, it is generally impossible to reformulate Push-Pull to obtain a centralized SGD-like update. The recursion of Push-Pull is
-$\displaystyle{\mathbf{x}}^{(k+1)}$ $\displaystyle=A{\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)},$ (8a)
-$\displaystyle{\mathbf{y}}^{(k+1)}$ $\displaystyle=B{\mathbf{y}}^{(k)}+{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)},$ (8b)
-where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the topology, $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$, ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$.
-Left-multiplying both sides of (8b) by $\mathds{1}^{\top}$, we achieve
-$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(k+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(k)}+\mathds{1}^{\top}{\mathbf{g}}^{(k+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k+1)},$ (9)
-where the last equality holds due to the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Next, we left-multiply both sides of (8a) by $\pi_{A}^{\top}$, which yields
-$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(k+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}$
-$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(k)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{k}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{k}.$
-Let $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$, and recall from (9) that $\mathds{1}^{\top}{\mathbf{y}}^{(k)}=\mathds{1}^{\top}{\mathbf{g}}^{(k)}$. Then, we have
-$\displaystyle\hat{x}^{(k+1)}=\hat{x}^{(k)}-c\alpha\bar{g}^{(k)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)},$ (10)
-where $c=n\pi_{A}^{\top}\pi_{B}$ is a constant and $\bar{g}^{(k)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(k)}$.
-We observe that the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}$ does not vanish but remains a large constant, see Figure 1.
-This non-vanishing error term makes achieving linear speedup in Push-Pull highly nontrivial, because in traditional linear speedup analysis, only the noise from $\|\bar{g}^{(k)}\|^{2}$ is the dominating factor.
-But here, both the squared error term $\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\|^{2}$ and the cross product $\left\langle\bar{g}^{(k)},\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}\right\rangle$ can introduce additional noise that may dominate the noise in $\bar{g}^{(k)}$.
-For certain specific topologies, the error term can be ignored or absorbed.
-For instance, the work in [5] explores a special case known as the $B$-ary tree Push-Pull (BTPP) and demonstrates the linear speedup result on this particular topology.
-In BTPP, $\pi_{A}=\pi_{B}=[1,0,\ldots,0]^{\top}$, leading to $\pi_{A}^{\top}(I-B_{\infty})=(0,-1,-1,\ldots,-1)$.
-By setting the loss function at the first node to zero, $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(k)}=-n\bar{g}^{(k)}$ is achieved, which simplifies the process of achieving linear speedup.
-Another case is when $A$ is doubly-stochastic; we have $\pi_{A}^{\top}(I-B_{\infty})=0$, which removes the error term.
-Unfortunately, for general topologies, the error term lacks exploitable structure.
-This usually leads to a significant overestimation of the noise term, which in turn obstructs the derivation of a linear speedup result.
-Therefore, we recognize the central challenge in the analysis of stochastic Push-Pull: How to estimate the noise level in error terms more accurately?
-We tackle this key question in the following section.
+FROST/PullDiag-GT is a Pull-Only method based on a row-stochastic matrix $A$. Its updates are
+$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (9a)
+$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=A{\mathbf{y}}^{(t)}+D_{t+1}^{-1}{\mathbf{g}}^{(t+1)}-D_{t}^{-1}{\mathbf{g}}^{(t)},$ (9b)
+where $D_{t}=\mathrm{Diag}(A^{t})$, ${\mathbf{g}}^{(t)}$ is the stacked stochastic gradient, and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying (9a) by $n^{-1}\pi_{A}^{\top}$ gives
+$n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\bar{g}^{(t)}+\epsilon^{(t)},$
+where $\epsilon^{(t)}=\alpha n^{-1}(\mathds{1}^{\top}-\pi_{A}^{\top}D_{t}^{-1}){\mathbf{g}}^{(t)}$ decays exponentially fast [39, Lemma 12].
+Setting $\hat{x}^{(t)}=n^{-1}\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ recovers the form (7).
+3.2 Push–Pull does not explicitly align with centralized SGD
+Unlike Push-Only and Pull-Only, the Push-Pull method cannot be perfectly aligned with centralized SGD.
+Although it can be written in the form of (7), it suffers from a non-vanishing error term.
+Recall the Push-Pull recursion (2).
+Left-multiplying (2b) by $\mathds{1}^{\top}$ yields
+$\displaystyle\mathds{1}^{\top}{\mathbf{y}}^{(t+1)}=\mathds{1}^{\top}{\mathbf{y}}^{(t)}+\mathds{1}^{\top}{\mathbf{g}}^{(t+1)}-\mathds{1}^{\top}{\mathbf{g}}^{(t)}=\mathds{1}^{\top}{\mathbf{g}}^{(t+1)},$ (10)
+where the last equality follows from the initialization ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. Left-multiplying step (2a) by $\pi_{A}^{\top}$, we obtain
+$\displaystyle\pi_{A}^{\top}{\mathbf{x}}^{(t+1)}$ $\displaystyle=\pi_{A}^{\top}A{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}B_{\infty}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$
+$\displaystyle=\pi_{A}^{\top}{\mathbf{x}}^{(t)}-\alpha\pi_{A}^{\top}\pi_{B}\mathds{1}^{\top}{\mathbf{y}}^{(t)}-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}.$
+Letting $\hat{x}^{(t)}=\pi_{A}^{\top}{\mathbf{x}}^{(t)}$ and using (10), we obtain
+$\displaystyle\hat{x}^{(t+1)}=\hat{x}^{(t)}-c\alpha\bar{g}^{(t)}+\epsilon^{(t)},\quad\text{where}\quad\epsilon^{(t)}=-\alpha\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)},$ (11)
+with $c=n\pi_{A}^{\top}\pi_{B}$ (see Definition (6)) and $\bar{g}^{(t)}=n^{-1}\mathds{1}^{\top}{\mathbf{g}}^{(t)}$.
+Unlike Push-DiGing (Example 1), where $\epsilon^{(t)}=0$, or FROST (Example 2), where $\epsilon^{(t)}$ decays exponentially, the error term $\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}$ in Push-Pull does not vanish but instead stabilizes at a non-negligible constant.
+As shown in Figure 1, this persistent error remains regardless of network size or topology, making the linear speedup analysis of Push-Pull substantially more challenging.
+Figure 1: With a fixed stepsize $\alpha$, the error term $\epsilon^{(t)}$ in Push-Pull does not vanish but rapidly stabilizes at a non-negligible constant.
+Results are shown for logistic regression on a synthetic dataset.
+4 A Novel Multiple-Step Descent Analysis Framework
+Establishing linear speedup for Push-Pull is challenging due to the presence of a non-vanishing error in (11).
+We begin by revisiting the classical single-step descent framework, highlighting why it suffices for Push-Only and Pull-Only methods but fails for Push-Pull, thereby motivating the development of a new multi-step descent analysis.
+4.1 Limitations of single-step descent analysis
+Following conventional single-step descent analysis, we establish the descent inequality based on update (11):
+$\displaystyle\|\nabla f(\hat{x}^{(t)})\|^{2}\hskip-2.84526pt\lesssim\hskip-2.84526pt\frac{f(\hat{x}^{(t)})\hskip-2.84526pt-\hskip-2.84526pt\mathbb{E}_{t}[f(\hat{x}^{(t+1)})]}{c\alpha}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\alpha L\|\pi_{A}^{\top}(I\hskip-2.84526pt-\hskip-2.84526ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}\hskip-2.84526pt+\hskip-2.84526pt\Delta_{\text{consent}}^{(t)}$ (12)
+where $\Delta_{\text{consent}}^{(t)}$ denotes consensus errors among local variables, typically of order $\mathcal{O}(\alpha^{2})$ and hence negligible when $\alpha$ is sufficiently small.
+We use $\lesssim$ to denote inequalities that hold up to absolute numerical constants.
+Taking full expectation and summing over $t=0,\dots,T-1$ yields
+$\displaystyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\lesssim\frac{f(\hat{x}^{(0)})\hskip-1.42262pt-\hskip-1.42262ptf(\hat{x}^{(T)})}{c\alpha T}\hskip-1.42262pt+\hskip-1.42262pt\frac{c\alpha L}{n}\sigma^{2}\hskip-1.42262pt\hskip-1.42262pt$ (13)
+$\displaystyle\quad+\frac{\alpha L}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I\hskip-1.42262pt-\hskip-1.42262ptB_{\infty}){\mathbf{y}}^{(t)}\|^{2}]+\frac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$].
+(14)
+Since $\tfrac{1}{T}\sum_{t=1}^{T-1}\mathbb{E}[\Delta^{(t)}_{\rm consent}$] is negligible for sufficiently small $\alpha$, the convergence rate is dominated by the two main terms in (13), provided the first term in (14) is either zero (Example 1) or decays exponentially (Example 2).
+With $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$, we achieve $\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]=\mathcal{O}(\sigma/\sqrt{nT})$, implying $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity and hence establish the linear speedup property.
+However, unlike Push-Only and Pull-Only, the term $\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty}){\mathbf{y}}^{(t)}\|^{2}]$ in Push-Pull persists (see Figure 1).
+The central challenge in establishing linear speedup is therefore to prove the following inequality:
+$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\pi_{A}^{\top}(I-B_{\infty})\mathbf{y}^{(t)}\|^{2}]\lesssim\frac{c}{n}\sigma^{2}+\text{negligible terms}.$ (15)
+If this inequality holds, the first term in (14) can be absorbed into the second term of (13).
+Choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ will yield the same linear speedup rate.
+Unfortunately, this cannot be proved with standard techniques.
+Recall that $B_{\infty}=\pi_{B}\mathds{1}_{n}^{\top}$ and define $\Delta_{y}^{(t)}:=(I-B_{\infty})\mathbf{y}^{(t)}$.
+Left-multiplying (2b) by $(I-B_{\infty})$ gives
+$\displaystyle\Delta_{y}^{(t+1)}=(B-B_{\infty})\Delta_{y}^{(t)}+(I-B_{\infty})\big({\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)}\big).$ (16)
+where ${\mathbf{g}}^{(t)}$ stacks the $n$ stochastic gradients $\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})$ (each with variance $\sigma^{2}$), we can therefore derive that $\mathbb{E}[\|\pi_{A}^{\top}\Delta_{y}^{(t)}\|^{2}]=\Theta(n\sigma^{2})$, which is far larger than the desired $c\sigma^{2}/n$.
+This motivates the need for a fundamentally different approach.
+4.2 A novel multi-step descent analysis
+In contrast to analyzing each iteration in isolation, we now introduce a novel $m$-step recursion that aggregates information across a block of $m$ iterations.
+This blockwise perspective reveals cancellation effects in the gradient-tracking recursion and enables the gradient noise to be controlled at the desired level.
+Specifically, for a positive integer $m$ and any index $k$ that is a multiple of $m$, the recursion (11) can be written as
+$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},$ (17)
+where $\Delta_{y}^{(t)}=(I-B_{\infty})\mathbf{y}^{(t)}$. Analogous to the derivation of the single-step descent inequality (12), we can establish the following $m$-step descent inequality:
+$\displaystyle\|\nabla f(\hat{x}^{(k)})\|^{2}\hskip-2.27621pt\lesssim\hskip-2.27621pt\frac{f(\hat{x}^{(k)})\hskip-2.27621pt-\hskip-2.27621pt\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]}{c\alpha m}\hskip-2.27621pt+\hskip-2.27621pt\frac{c\alpha L}{n}\sigma^{2}\hskip-2.27621pt+\hskip-2.27621pt\frac{\alpha L}{m}\|\pi_{A}^{\top}\hskip-2.27621pt\sum_{i=0}^{m-1}\hskip-2.84526pt\Delta_{y}^{(k+i)}\hskip-0.56905pt\|^{2}\hskip-2.27621pt+\hskip-2.27621pt\Delta_{\text{consent}}^{(k,m)}$ (18)
+where $\Delta_{\text{consent}}^{(k,m)}$ denotes consensus errors within the $m$-step block, which are typically $\mathcal{O}(\alpha^{2})$ and negligible for small $\alpha$.
+A formal statement of (18) is provided in Lemma 3.
+Let $T=mK$ denote the total number of iterations and define $[m^{*}]:=\{0,m,2m,\ldots,(K-1)m\}$.
+Taking the full expectation of (18) and summing over all $k\in[m^{*}]$ yields
+$\displaystyle\quad\min_{0\leq t\leq T-1}\mathbb{E}[\|\nabla f(\hat{x}^{(t)})\|^{2}]$ $\displaystyle\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\lesssim\frac{f(\hat{x}^{(0)})-f(\hat{x}^{(T)})}{c\alpha T}+\frac{c\alpha L}{n}\sigma^{2}$
+$\displaystyle\hskip-22.76228pt+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}]+\frac{\alpha L}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{\text{consent}}^{(k,m)}].$ (19)
+Similar to the single-step descent inequality (14), the central component in the $m$-step analysis is to control $\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}$.
+Summing the single-step recursion (16) from $i=0$ to $m-1$ yields
+$\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Delta_{y}^{(k)}+(B-B_{\infty})\sum_{i=0}^{m-2}\Delta_{y}^{(k+i)}+(I-B_{\infty})({\mathbf{g}}^{(k+m-1)}-{\mathbf{g}}^{(k)}).$
+This telescoping structure cancels all intermediate stochastic gradients ${\mathbf{g}}^{(k+i)}$ for $1\leq i\leq m-2$, leaving only boundary terms $\mathbf{g}^{(k)}$ and $\mathbf{g}^{(k+m-1)}$.
+Repeatedly applying this argument yields the equality
+$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\Big(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\Big)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (20)
+Unlike the single-step recursion (15), the $m$-step formulation cancels all interior noise, leaving only boundary differences ${\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}$, each attenuated by exponentially decaying factors $B^{m-j-1}-B_{\infty}$ (see Proposition 2).
+As a result, the block-level noise remains bounded at a constant level, independent of $m$.
+In particular, each boundary difference contributes gradient noise with variance $n\sigma^{2}$, but its effect is scaled by $\lambda_{B}^{m-j-1}$, yielding $\sum_{j=1}^{m-1}\lambda_{B}^{m-j-1}n\sigma^{2}\leq(n\sigma^{2})/(1-\lambda_{B})$ which is constant in $m$.
+Therefore,
+$\displaystyle\ \frac{1}{T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\pi_{A}^{\top}\Big(\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\Big)\|^{2}]$ (21)
+$\displaystyle\lesssim$ $\displaystyle\ \frac{n\sigma^{2}}{m}+\text{negligible terms}\lesssim\frac{\sigma^{2}}{n}+\text{negligible terms}.$
+The second inequality holds when $m$ is chosen sufficiently large [^1] In rigorous proofs, the choice of $m$ is constrained by additional factors and is therefore more involved than the simplified condition stated here, see more details in Section 5., e.g., $m\geq\text{Const}\cdot n^{2}$.
+Recalling that $T=mK$, substituting (21) into (18) and choosing $\alpha=\sqrt{n/(c^{2}\sigma^{2}T)}$ yield the desired linear speedup rate.
+This highlights the unique advantage of the multiple-step descent analysis: by aggregating errors across $m$ steps within a block, it reveals a telescoping cancellation in the gradient-tracking recursion, leaving only boundary terms that are exponentially decayed during iterations.
+Consequently, the effective gradient noise is controlled at the desired $\mathcal{O}(\sigma^{2}/n)$ level, enabling the first proof of linear speedup for Push-Pull on strongly connected digraphs.
 5 Achieving Linear Speedup for Push-Pull
-For simplicity, we denote $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$.
-Using the gradient tracking recursion (8b), we discover the following relationship about extra error terms: for any integer $k\geq 0,m\geq 1$:
-$\displaystyle\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}=\left(\sum_{j=0}^{m-1}(B^{j}-B_{\infty})\right)\Delta_{y}^{(k)}+\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-{\mathbf{g}}^{(k)}).$ (11)
-To understand this equation, we propose the following important supporting lemma:
-Lemma 3.
-For any $l\geq 0$ and any matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$.
-For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
-$\displaystyle\mathbb{E}_{k}[\|H\|_{F}^{2}]\leq n\sigma^{2}\sum_{i=0}^{l}\|M^{(i)}\|_{2}^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{l}M^{(i)}\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}].$
-Proof.
-This is a natural corollary of Assumption 3. ∎
-Now we turn to the noise estimation for $\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]$. A common analysis using Cauchy’s inequality typically provides
-$\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\leq m\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim m^{2}\sigma^{2}.$ (12)
-Here, we use $\lesssim$ to absorb matrix-related constants and focus on $\sigma^{2}$-related terms.
-Nevertheless, analysis using Equation (11) and Lemma 3 can be significantly better.
-Proposition 2 ensures that $\textstyle\sum_{j=1}^{m-1}\|B^{m-j-1}-B_{\infty}\|_{2}^{2}+\|\sum_{j=1}^{m-1}B^{m-j-1}-B_{\infty}\|_{2}^{2}$ is finite.
-Therefore, applying Lemma 3 to Equation (11) tells us
-$\displaystyle\mathbb{E}_{k}[\|\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|_{F}^{2}]\lesssim\mathbb{E}_{k}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+\sigma^{2}$ (13)
-$\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
-$\displaystyle\lesssim\sigma^{2}+\alpha^{2}L^{2}m^{2}\sigma^{2}.$
-Estimate from (13) sufficiently exploits the structure of gradient tracking, thus exhibiting ${\mathcal{O}}(m^{2})$ improvement compared to a naive estimate (12).
-To effectively leverage the result of  (13), we introduce an $m$-step descent lemma.
-It comes from applying the standard descent lemma to the following recursion:
-$\hat{x}^{(k+m)}=\hat{x}^{(k)}-c\alpha\sum_{i=0}^{m-1}\bar{g}^{(k+i)}-\alpha\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}.$ (14)
+In this section, we formalize the aforementioned arguments through supporting lemmas and rigorously establish the linear speedup property for Push-Pull.
+5.1 Supporting lemmas
 Lemma 3 (m-step Descent Lemma).
@@
 For any integers $k\geq 0$, we define $\hat{x}^{(k)}=\pi_{A}^{\top}{\mathbf{x}}^{(k)}$.
-When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for any integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$:
-$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (15)
-$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}$
-$\displaystyle\quad+\frac{c\alpha L^{2}}{2}\Delta_{1}^{(k)}+\frac{c\alpha L^{2}}{2}\Delta_{2}^{(k)}+\frac{3\alpha s_{B}^{2}}{cm}\|\Delta_{y}^{(k)}\|_{F}^{2},$
-where
-$\displaystyle\Delta_{1}^{(k)}=\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
-The remaining errors are estimated in the next lemma.
-Lemma 5 (Remaining Errors).
+When $\alpha\leq\frac{1}{6cmL}$, the following inequality holds for integers $k\geq 0,m\geq 6c^{-2}s_{B}^{2}$ (Proof is in Appendix B.1):
+$\displaystyle\quad\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{1}{m}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (22)
+$\displaystyle\leq\frac{4(f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})])}{\alpha cm}+\frac{8\alpha cL}{n}\sigma^{2}+\frac{8\alpha Lns_{B}^{2}}{cm}\sigma^{2}$ (23)
+$\displaystyle\quad+\frac{2L^{2}}{m}\Delta_{1}^{(k)}+\frac{2L^{2}}{m}\Delta_{2}^{(k)}+\frac{12s_{B}^{2}}{c^{2}m^{2}}\|\Delta_{y}^{(k)}\|_{F}^{2},$ (24)
+where $\overline{\nabla f}^{(k+i)}:=n^{-1}\mathbf{1}^{\top}\nabla f({\mathbf{x}}^{(k+i)})$ denotes the average of the local deterministic gradients evaluated at ${\mathbf{x}}^{(k+i)}$, and
+$\displaystyle\Delta_{1}^{(k)}:=\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-\hat{{\mathbf{x}}}^{(k)}\|_{F}^{2}],\quad\Delta_{2}^{(k)}:=\sum_{j=1}^{m}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$
+In (23), when $m\geq c^{-2}n^{2}s_{B}^{2}$, the noise terms associated with $\sigma^{2}$ are dominated solely by the linear speedup term $\tfrac{8\alpha cL}{n}\sigma^{2}$, which is consistent with our analysis in (21).
+The remaining error terms appearing in (24) are estimated in the next lemma.
+Lemma 4.
 Suppose Assumption 34 and Proposition 1 hold.
 With $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ defined above, when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\}$ and $m\geq C_{m}$, we have
-$\displaystyle\quad\frac{c\alpha L^{2}}{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle\leq\frac{\alpha c}{4}\sum_{t=0}^{mK-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{\alpha cm}{4}(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
-where we define $[m^{*}]:=\{0,m,\ldots,(K-1)m\}$, and $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants.
-$s_{B^{m}}$ is the measure defined for $B^{m}$, see (3).
-Generally, we have $s_{B^{m}}\leq s_{B}$, and $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}$ for large $m$.
-We are ready to state the main result of this paper.
+$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq\frac{1}{m}\sum_{t=0}^{T-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(t)}\|^{2}]+(\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}})\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{4m^{2}}.$
+Here, $C_{m}$ and $C_{\Delta,1}$ denote matrix-related constants (see (B.2)).
+The quantity $s_{B^{m}}$ represents the matrix measure of $B^{m}$, as defined in (5).
+In general, we have $s_{B^{m}}\leq s_{B}$, and moreover, $s_{B^{m}}$ converges to $\lVert I-B_{\infty}\rVert_{2}^{2}$ as $m\to\infty$ (Proof is in Appendix B.2).
+With the above lemmas, we are ready to state the main result of this paper.
 Theorem 1.
@@
 $\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (25)
-where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants. Furthermore, if we choose
+where $C_{m}$ and $C_{\Delta,1}$ are matrix-related constants defined in (B.2) and (53). Furthermore, if we choose
 $\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
 $\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
-we reach that
+it holds that
 $\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]$ $\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$ (26)
@@
 Proof.
-See Appendix B.3
+See Appendix B.3. ∎
 Comparison with existing analyses.
-Theorem 1 provides the first finite-time non-convex convergence result for stochastic Push-Pull, while previous analyses focus either on the strongly convex case [46, 40] or on variants of Push-Pull relying on specific topologies [5].
-The first term $\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ explicitly shows that Push-Pull enjoys a linear speedup on arbitrary topologies satisfying Proposition 2.
-Impact of $m$: In our analysis, we consider $m$-step gradient descent.
-The parameter $m$ effectively controls the terms related to $\sigma^{2}$ from additional errors, as illustrated by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (16).
-$m$ cannot be arbitrarily large because larger values of $m$ restrict the range of the learning rate, resulting in slower convergence and longer transient times in the analysis.
-In summary, the parameter $m$ allows for a tradeoff between noise level and transient time.
-Transient time. Our theorem is able to provide a coarse estimate for the transient time, which is
-$\max\{n^{2}c^{-4}s_{B}^{8},n^{7}c^{-4}s_{A}^{2}s_{B}^{3},nc^{-4}s_{B}^{4}\}.$
-The order of $n$ is overestimated because we frequently use $\|\pi_{A}\|\leq 1$ and $\|\pi_{B}\|\leq 1$, leading to a loss of $n$ when $A$ or $B$ is doubly-stochastic.
+Theorem 1 provides the first convergence guarantee for stochastic Push-Pull on nonconvex problems.
+In contrast, prior analyses focus either on the strongly convex setting [42, 25] or on Push-Pull variants restricted to specific network topologies [7].
+The leading term $\tfrac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}$ demonstrates a linear speedup with respect to the number of nodes $n$ for any topology satisfying Proposition 2.
+Consequently, Push-Pull requires $\mathcal{O}(1/(n\epsilon^{2}))$ iterations to achieve an $\epsilon$-accurate solution, with the complexity decreasing linearly as $n$ increases.
+Impact of $m$: In our analysis, we consider $m$-step gradient descent, where the parameter $m$ governs the noise-related terms, as reflected by the $\mathcal{O}(m^{-2}\sigma^{2})$ term in (25).
+However, $m$ cannot be taken arbitrarily large: increasing $m$ reduces the admissible range of the learning rate, thereby slowing convergence in the analysis.
+Overall, $m$ balances noise reduction against convergence rate.
+With an appropriate choice of $m$, one can suppress noise, achieve linear speedup, and still guarantee $\mathcal{O}(1/\sqrt{nT})$.
 6 Experiment
-In this section, we empirically investigate the linear speedup properties of the Push-Pull algorithm across diverse network topologies.
-For the stochastic gradient oracle, we ensure that each node operates on a local dataset, computing gradients from randomly sampled data points at each iteration.
-The experimental validation proceeds in two phases: first, we examine a nonconvex regression problem using synthetic data, followed by training a fully connected neural network on the heterogeneously distributed MNIST dataset.
-To further assess practical performance on established benchmarks, we extend our evaluation to ResNet-18 training on CIFAR-10.
-The distributed training experiments are conducted across multiple nodes ($n>1$) using three distinct network topologies: directed exponential graphs, grid graphs, and undirected random graphs.
-All results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup achieved through distributed optimization.
-Complete implementation specifications are provided in Appendix C. The source code and experimental scripts are publicly available in our GitHub repository.
+In this section, we empirically examine the linear speedup properties of the Push-Pull algorithm across a variety of network topologies.
+Each node in the stochastic gradient oracle operates on a local dataset, computing gradients from randomly sampled data points at every iteration.
+The experimental study consists of two phases: (i) nonconvex regression on synthetic data, and (ii) training a fully connected neural network on the heterogeneously distributed MNIST dataset.
+To further evaluate practical performance on standard benchmarks, we extend our experiments to training ResNet-18 on CIFAR-10.
+The distributed training is performed across multiple nodes ($n>1$) under different network topologies, including directed exponential graphs, grid graphs, and random graphs.
+For reference, all results are systematically compared against the single-node baseline ($n=1$) to quantify the speedup obtained through distributed optimization.
+Complete implementation details are provided in Appendix C, and the source code with experimental scripts is publicly available at our GitHub repository.
 6.1 Nonconvex Logistic Regression on Synthetic Dataset
-In the first experiment, the loss function at each node is a logistic regression function with nonconvex regularization terms (see (44) in Appendix C).
-We studied six different network topologies: the exponential, grid, and ring graphs are directed, while the random, geometric, and nearest neighbor graphs are undirected.
-The mixing matrices $A$ and $B$ are generated by a special procedure (see Appendix C.1), which produces row-stochastic or column-stochastic matrices, but not doubly-stochastic ones.
-We can observe clear linear speedup curves in Figure 2.
-For each topology, the noise level in gradients decreases proportionally to the square root of the number of nodes after the same number of communication rounds.
-All learning rates are set to $lr=0.005$.
-Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all 0.0050.005 0.0050.005.
+In the first experiment, the loss function at each node is logistic regression with nonconvex regularization terms.
+We evaluate six network topologies: exponential, grid, and ring graphs (directed), and random, geometric, and nearest-neighbor graphs (undirected).
+The mixing matrices $A$ and $B$ are constructed using a specialized procedure, which yields row-stochastic or column-stochastic matrices, but not doubly stochastic ones.
+Details are provided in Appendix C.1 and Appendix C.2.
+As shown in Figure 2, the results exhibit clear linear speedup.
+Across all topologies, the gradient noise level decreases proportionally to the square root of the number of nodes, given the same number of communication rounds.
+The learning rate is fixed at $lr=0.005$ for all experiments.
+Figure 2: Linear speedup performance over various topologies on synthetic data, the learning rates are all $0.005$.
 6.2 Training Neural Network for MNIST Classification
 In the second experiment, we trained a three-layer fully connected neural network for MNIST classification.
-We used the same topologies as before: directed exponential, grid, and ring graphs, and undirected random, geometric, and nearest neighbor graphs.
-Data is distributed heterogeneously using a Directlet distribution, see Appendix C.3.
-Figure 3 shows that Push-Pull is able to achieve linear speedup on both directed and undirected topologies.
+The same network topologies as in the synthetic dataset experiment were used.
+The MNIST data are distributed heterogeneously according to a Dirichlet distribution (see Appendix C.3).
+As shown in Figure 3, Push-Pull achieves linear speedup on both directed and undirected topologies.
 Figure 3: Linear speedup performance over various topologies, the learning rates are the same over the same topology.
 6.3 Training Neural Network for CIFAR10 Image Classification
-In the third experiment, we trained Resnet18 for CIFAR10 classification across different numbers of nodes and various topologies.
-We used three topologies mentioned before: directed exponential, grid graphs, and undirected random graphs.
-Here $n=1$ means training a single Resnet18 on a single-node without distributed optimization.
-Figure 4 demonstrates the performance of our Push-Pull method.
-We observe that Push-Pull’s accuracy is initially lower compared to single-node optimization, which can be attributed to the overhead of the influence of topology.
-Nevertheless, after a sufficient number of epochs, Push-Pull achieves superior accuracy, leveraging its inherent speed advantages.
+In the third experiment, we trained ResNet-18 for CIFAR-10 classification across varying numbers of nodes and network topologies.
+We considered three representative topologies: directed exponential graphs, grid graphs, and undirected random graphs.
+For reference, $n=1$ corresponds to training ResNet-18 on a single node without distributed optimization.
+Figure 4 illustrates the performance of the Push-Pull method.
+We observe that its accuracy is initially lower than that of single-node training, due to the overhead introduced by network topology.
+However, after a sufficient number of epochs, Push-Pull surpasses the single-node baseline, benefiting from its inherent speedup advantages.
 Figure 4: The test accuracy across different numbers of nodes and various topologies.
-All experiments were conducted using a learning rate of 0.0050.005 0.0050.005.
-The lower three figures specifically depict the accuracy after 100100 100100 epochs.
+All experiments were conducted using a learning rate of $0.005$.
+The lower three figures specifically depict the accuracy after $100$ epochs.
 7 Conclusion
@@
 Numerical experiments have corroborated our theoretical findings.
-Appendix A Rolling-Sum Lemma
+Appendix A Linear Algebra Inequalities
 Lemma 5 (Rolling Sum Lemma).
@@
 $\displaystyle\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}=\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}\|A^{i}-A_{\infty}\|_{2}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}\|A^{i}-A_{\infty}\|_{2}$ (30)
-[…]
-[…]
+$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{\min\{\lfloor p\rfloor,k\}}M_{A}+\sum_{i=\min\{\lfloor p\rfloor,k\}+1}^{k}M_{A}\beta_{A}^{i-p}$
+$\displaystyle\leq$ $\displaystyle M_{A}\cdot(1+\min\{\lfloor p\rfloor,k\})+M_{A}\cdot\frac{1}{1-\beta_{A}}\beta_{A}^{\min\{\lfloor p\rfloor,k\}+1-p}.$
 If $p=0$, (30) is simplified to $\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\leq M_{A}\cdot\frac{1}{1-\beta_{A}}$ and (29) is naturally satisfied.
@@
 $\displaystyle\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}=\|\sum_{i=0}^{k}a_{k-i}\cdot a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}$ (31)
-[…]
+$\displaystyle\leq$ $\displaystyle\sum_{i=0}^{k}a_{k-i}\|a_{k-i}^{-1}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq\sum_{i=0}^{k}a_{k-i}^{-1}\|A^{k-i}-A_{\infty}\|_{2}^{2}\|\Delta^{(i)}\|_{F}^{2}.$
 By choosing $a_{k-i}=(\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2})^{-1}\|A^{k-i}-A_{\infty}\|_{2}$ in (31), we obtain that
@@
 $\displaystyle\sum_{k=0}^{T}\|\sum_{i=0}^{k}(A^{k-i}-A_{\infty})\Delta^{(i)}\|_{F}^{2}\leq s_{A}\sum_{k=0}^{T}\sum_{i=0}^{k}\|A^{k-i}-A_{\infty}\|_{2}\|\Delta^{(i)}\|_{F}^{2}$
-[…]
+$\displaystyle\leq$ $\displaystyle s_{A}\sum_{i=0}^{T}(\sum_{k=i}^{T}\|A^{k-i}-A_{\infty}\|_{2})\|\Delta^{(i)}\|_{F}^{2}\leq s_{A}^{2}\sum_{i=0}^{T}\|\Delta^{(i)}\|_{F}^{2},$
 which finishes the proof of this lemma. ∎
+Lemma 6.
+Given $l\geq 0$ and matrices $M^{(0)},M^{(1)},\ldots,M^{(l)}\in\mathbb{R}^{n\times n}$.
+For any $k\geq 0$, If $H=\sum_{i=0}^{l}M^{(i)}{\mathbf{g}}^{(k+i)}$ is a linear combination of stochastic gradients, then
+$\displaystyle\mathbb{E}_{k}[\|H\|_{F}^{2}]\leq n\sigma^{2}\sum_{i=0}^{l}\|M^{(i)}\|_{2}^{2}+\mathbb{E}_{k}[\|\sum_{i=0}^{l}M^{(i)}\nabla f({\mathbf{x}}^{(k+i)})\|_{F}^{2}].$
+Proof.
+This is a direct corollary from the unbiasedness and linearly independent gradients in Assumption 3. ∎
 Appendix B Proofs of Main Theorem
 B.1 Proof of Lemma 3
-For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x^{(k)}_{i})$. According to Assumption 3, we have
+For simplicity, we define $\Delta_{y}^{(k)}:=(I-B_{\infty}){\mathbf{y}}^{(k)}$ and $\overline{\nabla f}^{(k)}:=n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})$. According to Assumption 3, we have
 $\mathbb{E}_{k}[\bar{g}^{(k+i)}]=\mathbb{E}_{k}[\mathbb{E}_{k+i}[\bar{g}^{(k+i)}]]=\mathbb{E}_{k}[\overline{\nabla f}^{(k+i)}],\quad\forall i\geq 0.$
@@
 $\displaystyle\quad+c^{2}\alpha^{2}L\|\sum_{i=0}^{m-1}\bar{g}^{(k+i)}\|^{2}+\alpha^{2}L\|\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}\|^{2}.$
-By taking expectation conditioned on ${\mathcal{F}}_{k}$, we obtain that
+Take expectation conditioned on ${\mathcal{F}}_{k}$ and we obtain that
 $\displaystyle\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]-f(\hat{x}^{(k)})\leq-c\alpha\sum_{i=0}^{m-1}\mathbb{E}_{k}\left[\left\langle\overline{\nabla f}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (33)
@@
 $\mathbb{E}_{k}\left[\left\langle\bar{g}^{(k+i)}-\overline{\nabla f}^{(k+i)},\bar{g}^{(k+j)}-\overline{\nabla f}^{(k+j)}\right\rangle\right]=0.$
-Plug these estimations into (24), we obtain that
+Plug the estimation for the inner product into (33), we obtain that
 $\displaystyle\frac{\alpha cm}{2}\mathbb{E}_{k}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]\leq\mathbb{E}_{k}[f(\hat{x}^{(k)})-f(\hat{x}^{(k+m)})]$ (34)
@@
 $\displaystyle=\sum_{j=1}^{m-1}\mathbb{E}_{k}[\|(B^{m-j-1}-B_{\infty})({\mathbf{g}}^{(k+j)}-\nabla f({\mathbf{x}}^{(k+j)}))\|_{F}^{2}]$
-$\displaystyle\quad+\mathbb{E}_{k}[\|\left(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})\right)({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
+$\displaystyle\quad+\mathbb{E}_{k}[\|(\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty}))({\mathbf{g}}^{(k)}-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
 $\displaystyle\quad+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
@@
 $\displaystyle\leq 2ns_{B}^{2}\sigma^{2}+\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-The inner product can produce a similar term:
-$\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)}]=\mathbb{E}_{k}[H_{m,k}].$
-Using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
-$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]=-\alpha\left\langle\mathbb{E}_{k}[H_{m,k}],\nabla f(\hat{x}^{(k)})\right\rangle$ (28)
+Similarly, using the fact that $\hat{x}^{(k)}$ is known under filtration ${\mathcal{F}}_{k}$, we have
+$\displaystyle\quad-\alpha\mathbb{E}_{k}\left[\left\langle\pi_{A}^{\top}\sum_{i=0}^{m-1}\Delta_{y}^{(k+i)},\nabla f(\hat{x}^{(k)})\right\rangle\right]$ (37)
 $\displaystyle=-\alpha\left\langle\mathbb{E}_{k}[\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))],\nabla f(\hat{x}^{(k)})\right\rangle$
 $\displaystyle\quad+\alpha\left\langle\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)},\nabla f(\hat{x}^{(k)})\right\rangle$
-[…]
-$\displaystyle\quad+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
+$\displaystyle\overset{\text{Cauchy}}{\leq}\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\pi_{A}^{\top}\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|^{2}]$ (38)
+$\displaystyle\quad+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{2\alpha}{cm}\|\pi_{A}^{\top}\left(\sum_{i=0}^{m-1}(B^{m-i-1}-B_{\infty})\right)\Delta_{y}^{(k)}\|_{F}^{2}+\frac{\alpha cm}{8}\|\nabla f(\hat{x}^{(k)})\|^{2}$
 $\displaystyle\leq\frac{2\alpha}{cm}\mathbb{E}_{k}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]+\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}$
@@
 $\displaystyle+\frac{2\alpha^{2}c^{2}mL}{n}\sigma^{2}+2\alpha^{2}Lns_{B}^{2}\sigma^{2}+\frac{3\alpha s_{B}^{2}}{cm}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}]$
-We can further apply Cauchy’s inequality on the last term in (31), i.e.
+$\displaystyle+\frac{3\alpha}{cm}\mathbb{E}[\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}].$
+We can further apply Cauchy’s inequality on the last term in (41), i.e.,
 $\displaystyle\quad\|\sum_{j=1}^{m-1}(B^{m-j-1}-B_{\infty})(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)}))\|_{F}^{2}$
@@
 $\displaystyle\leq s_{B}\sum_{j=1}^{m-1}\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}\leq s_{B}L^{2}\sum_{j=1}^{m-1}\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}.$
-Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (31) can be further simplified to
+Finally, when $m\geq 6c^{-2}s_{B}$, we have $\frac{3\alpha s_{B}L^{2}}{cm}\leq\frac{c\alpha L^{2}}{2}$, and (41) can be simplified to
 $\displaystyle\quad\frac{\alpha cm}{4}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}\sum_{i=0}^{m-1}\mathbb{E}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$
@@
 $\displaystyle+\frac{c\alpha L^{2}}{2}\sum_{j=1}^{k}\mathbb{E}_{k}[\|{\mathbf{x}}^{(k+j)}-{\mathbf{x}}^{(k)}\|_{F}^{2}],$
-which finishes the proof.
+By dividing $\frac{\alpha cm}{4}$ on both sides of this inequality, we finish the proof.
 B.2 Proof of Lemma 4
-In this section, we consider how to measure the difference of ${\mathbf{x}}$. To be more specific, we mainly consider the relationships between the following 44 44 terms:
-$\Delta_{1}^{(k)},\Delta_{2}^{(k)},\mathbb{E}_{k}[\|\Delta_{x}^{(k)}\|_{F}^{2}],\Delta_{y}^{(k)},\quad k\in[m^{*}].$
-Lemma 7 will give upper bounds of $\Delta_{1}^{(k)}$, $\Delta_{2}^{(k)}$ using other terms, while Lemma
-Lemma 7.
-Suppose Proposition 2 and Assumption 34 hold.
-When $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold for any positive integers $k$ and $m$.
+We first provide the complete definition of the constant $C_{m}$ used in the statement.
+We define $C_{m}$ as the maximum of all lower bounds imposed on $m$ to ensure the validity of subsequent inequalities and simplifications:
+$\displaystyle C_{m}=\max\Big\{$ $\displaystyle n^{-1}s_{A}^{2},20c^{-2}ns_{A}s_{A^{m}},\;c^{-2}n^{2}s_{B}^{2},30c^{-2}s_{B}^{2}s_{B^{m}}$
+$\displaystyle c^{-2}n^{2}(8s_{A}^{2}s_{B}+2ns_{B}+2s_{A}s_{B}+\sqrt{n}s_{B}),20c^{-2}n^{3}s_{A}s_{A^{m}}s_{B}s_{B^{m}}\Big\}.$ (42)
+We define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1},j\geq 1$ and the following constants:
+$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}\leq 8s_{A}^{2}s_{B}+2ns_{B},$
+$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}\leq 2s_{A}s_{B}+\sqrt{n}s_{B}.$
+For all arguments used in this section, we suppose Proposition 2 and Assumption 34 hold.
+We first declare that, when $\alpha\leq\frac{1}{6s_{B}L}$, $m\geq\max\{c^{-2}n^{2}(N_{7}+N_{8}^{2}),n^{-1}s_{A}^{2},n^{-1}s_{A},c^{-2}n^{2}s_{A}^{2}\}$, the following inequalities hold $\forall k,m\geq 1$.
 $\displaystyle\Delta_{1}^{(k)}$ $\displaystyle\leq 40mn\|\Delta_{x}^{(k)}\|_{F}^{2}+40\alpha^{2}ms_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+20c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$ (43)
@@
 $\displaystyle\quad+20c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+j)}\|^{2}],$
-Proof.
-Define $\delta_{B}^{(0)}=I-B_{\infty}$, $\delta_{B}^{(j)}=B^{j}-B^{j-1}$, when $j\geq 1$. We have
+Note that
 $\displaystyle\quad{\mathbf{x}}^{(k+i)}-\hat{{\mathbf{x}}}^{(k)}-(A^{i}-A_{\infty})\Delta_{x}^{(k)}=-\alpha\sum_{j=0}^{i-1}A^{i-j-1}\Delta_{y}^{(k+j)}-\alpha\sum_{j=0}^{i-1}A^{i-j-1}B_{\infty}{\mathbf{y}}^{(k+j)}$
@@
 $\displaystyle\quad+5c^{2}\alpha^{2}m^{2}\sum_{j=0}^{m-1}\|\overline{\nabla f}^{(k+j)}\|^{2}+5c^{2}\alpha^{2}m^{2}n^{-1}\sigma^{2}$
-$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+ms_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
+$\displaystyle\leq 5s_{A}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}s_{B}+mns_{B})\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
 $\displaystyle\quad+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
 where in the first inequality, the $\sigma^{2}$ terms are separated using Lemma 6, and the second inequality comes from rolling sum.
-$\displaystyle N_{7}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}^{2}\leq 8s_{A}^{2}s_{B}+2n\|\pi_{A}\|^{2}s_{B}$
-$\displaystyle N_{8}$ $\displaystyle:=\sum_{t=0}^{\infty}\|\sum_{j=0}^{t}A^{t-j}\delta_{B}^{(j)}\|_{2}\leq 2s_{A}s_{B}+\sqrt{n}\|\pi_{A}\|s_{B}$
-For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$. This gives
-$\displaystyle\quad\Delta_{2}^{(k)}\leq 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}$
-$\displaystyle\quad+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
+For the decomposition of ${\mathbf{x}}^{(k+i)}-{\mathbf{x}}^{(k)}$, the only difference is that we replace $(A^{i}-A_{\infty})\Delta_{x}^{(k)}$ with $(A^{i}-I)\Delta_{x}^{(k)}$.
+This gives
+$\displaystyle\quad\Delta_{2}^{(k)}\leq$ $\displaystyle 5\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\|\Delta_{x}^{(k)}\|_{F}^{2}+10\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$
+$\displaystyle+5\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+5\alpha^{2}N_{8}^{2}L^{2}\Delta_{2}^{(k)}$
+$\displaystyle+5\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
 Easy to prove that $\sum_{i=1}^{m}\|A^{i}-I\|_{2}^{2}\leq 2s_{A}+2mn$. When $\alpha\leq\frac{1}{5N_{8}L}$, we have
-$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+m)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (34)
+$\displaystyle\Delta_{2}^{(k)}$ $\displaystyle\leq 20(s_{A}+mn)\|\Delta_{x}^{(k)}\|_{F}^{2}+20\alpha^{2}(s_{A}^{2}+mn)s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}$ (45)
 $\displaystyle\quad+10\alpha^{2}m(N_{7}+N_{8}^{2}+c^{2}mn^{-2})n\sigma^{2}+10\alpha^{2}(n^{2}s_{A}^{2}+c^{2}m^{2})\|\overline{\nabla f}^{(k+j)}\|^{2}.$
@@
 $\displaystyle\Delta_{y}^{(k+m)}=(B^{m}-B_{\infty})\Delta_{y}^{(k)}+\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$ (46)
-Define
-$\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)}).$
-Note that it can be rewritten as
+We define $\Delta_{3}^{(k)}=\sum_{i=0}^{m-1}(B^{m-1-i}-B_{\infty})({\mathbf{g}}^{(k+i+1)}-{\mathbf{g}}^{(k+i)})$, which can be written as
 $\displaystyle\Delta_{3}^{(k)}$ $\displaystyle=(I-B_{\infty}){\mathbf{g}}^{(k+m)}+(B-I){\mathbf{g}}^{(k+m-1}+(B^{2}-B){\mathbf{g}}^{(k+m-2)}$
 $\displaystyle\quad+\ldots+(B^{m-1}-B^{m-2}){\mathbf{g}}^{(k+1)}-(B^{m-1}-B_{\infty}){\mathbf{g}}^{(k)},$
-according to Lemma 3, we have
-$\displaystyle\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$
-where
+according to Lemma 6, we have $\mathbb{E}_{k}[\|\Delta_{3}^{(k)}\|_{F}^{2}]\leq n\sigma^{2}q_{B}+h_{m,k}$, where
 $\displaystyle h_{m,k}$ $\displaystyle:=\sum_{j=1}^{m}\delta_{B}^{(m-j)}\nabla f({\mathbf{x}}^{(k+j)})-(B^{m-1}-B_{\infty})\nabla f({\mathbf{x}}^{(k)}),$
@@
 $\displaystyle h_{m,k}$ $\displaystyle=\sum_{j=1}^{m}\delta_{B}^{(m-j)}(\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})).$
-By applying the rolling sum lemma, we have
+By applying the Jensen’s inequality, we have
 $\displaystyle\mathbb{E}_{k}[\|h_{m,k}\|_{F}^{2}]$ $\displaystyle\leq\sum_{t=0}^{m}\|\delta_{B}^{(t)}\|_{2}^{2}\sum_{j=1}^{m}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
@@
 $\displaystyle\sum_{k\in[m^{*}]}\|\Delta_{y}^{(k)}\|_{F}^{2}$ $\displaystyle\leq 7s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+6s_{B^{m}}nL\Delta+4s_{B}s_{B^{m}}^{2}L^{2}\sum_{k\in[m^{*}]}\Delta_{2}^{(k)}.$
-Plug in (33), we have
-$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}$
-$\displaystyle\,\,+6s_{B^{m}}nL\Delta+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
+Replace the $\Delta_{2}$ term with the estimate in (44), we reach that
+$\displaystyle(1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2})\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq$ $\displaystyle(7s_{B}s_{B^{m}}^{2}+80c^{2}m^{2}n^{-1}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2})Kn\sigma^{2}+6s_{B^{m}}nL\Delta$
+$\displaystyle+160s_{B}s_{B^{m}}^{2}L^{2}mn\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+80c^{2}m^{2}s_{B}s_{B^{m}}^{2}\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$
 When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}s_{B}s_{B^{m}}$, we have $1-160\alpha^{2}mL^{2}s_{B}^{2}s_{B^{m}}^{2}\geq 0.5$.
@@
 $\displaystyle\Delta_{x}^{(k+m)}=(A^{m}-A_{\infty})\Delta_{x}^{(k)}-\alpha\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}.$ (49)
-Define
-$\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)},$
-then we can iterate (38) and obtain that
-$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(k)}.$
-To further analyze $\Delta_{4}^{(k)}$, we expand it as
+For $k\in[m^{*}]$, define $\Delta_{4}^{(k)}:=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty}){\mathbf{y}}^{(k+i)}$, then we can iterate (49):
+$\Delta_{x}^{(k+m)}=-\alpha\sum_{j=0}^{k/m}(A^{k-jm}-A_{\infty})\Delta_{4}^{(jm)},\quad k\in[m^{*}].$
+Using the rolling sum lemma, we know that for $k\in[m^{*}]$,
+$\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]].$
+To further analyze $\mathbb{E}_{k}[\Delta_{4}^{(k)}]$, $k\in[m^{*}]$, we expand it as
 $\displaystyle\Delta_{4}^{(k)}$ $\displaystyle=\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\Delta_{y}^{(k+i)}+n\sum_{i=0}^{m-1}(A^{m-i}-A_{\infty})\pi_{B}\bar{g}^{(k+i)}$
@@
 This gives
-$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
+$\displaystyle\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+3ns_{A}\sigma^{2}$
 $\displaystyle\quad+6s_{A}s_{B}\sum_{j=i}^{m-1}\mathbb{E}_{k}[\|\nabla f({\mathbf{x}}^{(k+j)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]+12s_{A}s_{B}n\sigma^{2}$
-$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]$
-$\displaystyle\quad+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
-Using the rolling sum lemma, we have
+$\displaystyle\leq 3s_{A}s_{B}\|\Delta_{y}^{(k)}\|_{F}^{2}+3n^{2}s_{A}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|_{F}^{2}]+6s_{A}s_{B}L^{2}\Delta_{2}^{(k)}+15s_{A}s_{B}n\sigma^{2}.$
+Plug in our estimate on $\Delta_{4}^{(k)}$, we have
 $\displaystyle\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]]$
@@
 Remember that $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ share the same upper bound, we utilize (52), (51) and obtain that
-$\displaystyle\frac{c\alpha L^{2}}{2}\quad\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
-$\displaystyle\leq\frac{\alpha cm}{4}C_{\sigma,1}K\sigma^{2}+\frac{c\alpha}{4}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{c\alpha L\Delta}{4m},$
+$\displaystyle\quad\frac{2L^{2}}{m}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{12s_{B}^{2}}{c^{2}m^{2}}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
+$\displaystyle\leq C_{\sigma,1}K\sigma^{2}+\frac{1}{m}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{L\Delta}{m^{2}},$
 where
@@
 When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq c^{-2}30s_{B}^{1.5}$, we have
-$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
-This finishes the proof.
+$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$ (53)
+This finishes the proof of Lemma 4. Finally, we leave a brief discussion on $s_{B^{m}}$. Recall that
+$s_{B^{m}}=\max\{\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}\},$
+when $m$ increases, both $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}$ and $\sum_{i=1}^{\infty}\|B^{im}-B_{\infty}\|_{2}^{2}$ converge to zero exponentially fast, and $\|I-B_{\infty}\|_{2}\geq 1$.
+Thus, $s_{B^{m}}$ converges to $\|I-B_{\infty}\|_{2}^{2}$ exponentially fast as $m$ grows.
 B.3 Proof of Theorem 1
@@
 $\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$
-$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)$
+$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right).$
 Appendix C Experiment Details
@@
 Subsequently, based on the specific network topology definition, we set $W_{i,j}=1$ if there is a directed edge from node $i$ to node $j$.
-After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from 11 11 to 99 99 to all entries where $W_{i,j}=1$.
+After constructing this initial 0-1 matrix $W$, we then randomly assign integer weights from $1$ to $9$ to all entries where $W_{i,j}=1$.
 Finally, we perform either row normalization or column normalization on this weighted matrix to obtain a row-stochastic matrix $A$ or a column-stochastic matrix $B$, respectively.
 C.2 Synthetic data
-For experiments conducted on the synthetic dataset, the objective function is given by:
+For experiments conducted on the synthetic dataset, we minimize the objective:
 $\displaystyle f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad x\in\mathbb{R}^{d},$ (56)
@@
 $f_{i}(x)=\underbrace{\frac{1}{L_{i}}\sum_{l=1}^{L_{i}}\ln\left(1+\exp(-y_{i,l}h_{i,l}^{\top}x)\right)}_{\text{Logistic Loss}}+\underbrace{\rho\sum_{j=1}^{d}\frac{[x]_{j}^{2}}{1+[x]_{j}^{2}}}_{\text{Non-convex Regularization}},\quad L_{i}=L_{\text{Local}}/n.$
-Data Synthesis Process:
-  1. 1.
-    Global Data Generation:
-    • Generate optimal parameters: $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$
-    • Create global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with $h_{l,j}\sim\mathcal{N}(0,1)$
-    • Compute labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ via randomized thresholding:
+Data Synthesis Process.
+We generate the synthetic dataset through the following three-step procedure.
+(1) Global Data Generation: We first generate optimal parameters $x_{\text{opt}}\sim\mathcal{N}(0,I_{d})$ and create a global feature matrix $H\in\mathbb{R}^{L_{\text{total}}\times d}$ with entries $h_{l,j}\sim\mathcal{N}(0,1)$.
+Labels $Y\in\mathbb{R}^{L_{\text{total}}}$ with $y_{l}\in\{-1,+1\}$ are computed via randomized thresholding:
 $y_{l}=\begin{cases}1&\text{if }1/z_{l}>1+\exp(-h_{l}^{\top}x_{\text{opt}})\\ -1&\text{otherwise}\end{cases},\quad z_{l}\sim\mathcal{U}(0,1)$
-  2. 2.
-    Data Distribution:
-    • Partition $H$ and $L$ across $n$ nodes using:
-      $H^{(i)}=H[iM:(i+1)M,\;:\;],\;\quad Y^{(i)}=Y[iM:(i+1)M],\;\quad M=L_{\text{total}}/n$
-      where $L_{\text{total}}$ must be divisible by $n$
-  3. 3.
-    Local Initial Points:
-    $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d}),\quad\sigma_{h}=10$
-Gradient Computation: At each iteration, each node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$.
-The gradient computation consists of two components:
+(2) Data Distribution: We partition $H$ and $Y$ equally across $n$ nodes using $H^{(i)}=H[iM:(i+1)M,:]$ and $Y^{(i)}=Y[iM:(i+1)M]$ where $M=L_{\text{total}}/n$, requiring $L_{\text{total}}$ to be divisible by $n$.
+(3) Local Initial Points: Each node $i$ starts from $x_{i}^{\star}=x_{\text{opt}}+\varepsilon_{i}$ where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{h}^{2}I_{d})$ with $\sigma_{h}=10$.
+Gradient Computation. At each iteration, node $i$ independently computes its stochastic gradient by randomly sampling a minibatch of size $B$ from its local dataset of size $L_{i}=L_{\text{total}}/n$:
 $\nabla f_{i,B}(x)=\underbrace{-\frac{1}{B}\sum_{b=1}^{B}\frac{y_{i,b}h_{i,b}}{1+\exp(y_{i,b}h_{i,b}^{\top}x)}}_{\text{Logistic Loss Gradient}}+\underbrace{\rho\sum_{j=1}^{d}\frac{2[x]_{j}}{(1+[x]_{j}^{2})^{2}}}_{\text{Non-convex Regularization}}$
-where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly from the $L_{i}$ local samples without replacement.
-Notice that the gradient on each node $i$ is computed on the local parameter $x_{i}\in\mathbb{R}^{d}$.
-Implementation Details:
-  • Global dataset size $L_{\text{total}}=204800$, batch size $B=200$
-  • Dimension $d=10$, regularization $\rho=0.01$
-  • Node configurations: $n\in\{4,6,8,9,12,16,18,24,25\}$, and the least common multiple (LCM) is $3,600$.
-Notice that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
-Evaluation Metric: We track $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ as gradient norm, where
-$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\;M=L_{\text{Total}}/n.$
-The experimental results shown in our plots represent the average performance across 2020 2020 independent repetitions, where each trial is conducted with the fixed random $\text{seed}=42$ for reproducibility.
-C.3 Neural Network Experiment
-We first employ three-layer fully connected neural networks for MNIST classification.
-Data Distribution: For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector
-$\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$
-where $\mathbf{1}^{10}$ is a 1010 1010-dimensional vector of ones.
-Each $q_{k,c}$ represents the target proportion of samples from class $c$ that node $k$ should ideally hold.
-By definition of the Dirichlet distribution, $q_{k,c}\geq 0,\forall c$ and $\sum_{c=0}^{9}q_{k,c}=1,\forall k$.
-Denote $N_{k,c}^{\text{target}}=q_{k,c}\cdot\frac{M}{n}$ as the target number of samples of class $c$ to be assigned to node $k$.
-Ignoring the slight deviations caused by non-divisibility, the number of samples assigned to each node $k$ is approximately $60,000/n$.
-Here we set $\alpha=0.9$ to get the heterogeneous data and the Figure 5 shows the distribution of samples with 1616 1616 nodes.
-The Figure 2 shows that the Push Pull achieves linear speedup on synthetic data.
-Figure 5: Comparison of uniform (left) and heterogeneous (right) data distributions across 1616 1616 nodes ($\alpha=0.9$).
-Evaluation Metric: After each batch computation, we first average the gradients obtained from the $n$ neural networks.
-Then we compute the norm of this averaged gradient.
-This norm is subsequently normalized by the number of parameters P of a single neural network.
-In our experiments, $\textbf{P}=1,863,690$.
-$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{ij}}\right\|_{2}/\sqrt{\textbf{P}},\text{ the gradient }\mathbf{g}_{i,B_{i}}\text{ is computed on the $\displaystyle j$-th batch of node $\displaystyle i$}$ (45)
-The Figure 3 shows that the Push Pull achieves linear speedup on both directed and undirected topologies.
-Finally, we trained Resnet18 models on the CIFAR10 dataset to evaluate the performance of our Push-Pull method on a benchmark computer vision task.
-This involved conducting training across different numbers of nodes ($n>1$) utilizing the three previously described network topologies: directed exponential, grid graphs, and undirected random graphs.
-The performance was compared to single-node training ($n=1$) without distributed optimization.
-(see Figure 4).
-Implementation Details:
-  • Learning rate $lr=0.005$ on each topology and each number of node.
-  • Batch size $bs=128$ on each topology and each number of node.
+where the minibatch $\{h_{i,b},y_{i,b}\}_{b=1}^{B}$ is drawn uniformly without replacement from the $L_{i}$ local samples.
+The gradient at each node $i$ is computed on its local parameter $x_{i}\in\mathbb{R}^{d}$.
+Implementation Details.
+We use global dataset size $L_{\text{total}}=204800$ with batch size $B=200$, dimension $d=10$, and regularization parameter $\rho=0.01$.
+Node configurations include $n\in\{4,6,8,9,12,16,18,24,25\}$ with least common multiple (LCM) of $3,600$.
+Note that $L_{\text{total}}=1,440,000=3,600\cdot 200\cdot 2=\text{LCM}\cdot B\cdot 2$.
+Evaluation Metric. We track the gradient norm $\|n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})\|$ where
+$\displaystyle n^{-1}\mathds{1}^{\top}\nabla f({\mathbf{x}}^{(k)})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,M}(\bm{x}_{i}^{(k)}),\quad M=L_{\text{Total}}/n.$
+Results represent averages over 20 independent trials with fixed random seed 42 for reproducibility.
+C.3 Neural Network Experiments
+We evaluate our Push-Pull method on two benchmark tasks: MNIST classification using three-layer fully connected networks and CIFAR-10 classification using ResNet-18.
+Data Distribution.
+To simulate heterogeneous data distributions across nodes, we employ Dirichlet sampling.
+For each node $k\in\{1,\dots,n\}$, we sample a class proportion vector $\mathbf{q}^{k}=(q_{k,0},q_{k,1},\dots,q_{k,9})\sim\text{Dirichlet}(\alpha\cdot\mathbf{1}^{10})$, where $\mathbf{1}^{10}$ is a 10-dimensional vector of ones.
+Each $q_{k,c}$ represents the proportion of samples from class $c$ assigned to node $k$, satisfying $q_{k,c}\geq 0$ and $\sum_{c=0}^{9}q_{k,c}=1$.
+The target number of samples of class $c$ for node $k$ is $N_{k,c}^{\text{target}}=q_{k,c}\cdot M/n$, resulting in approximately $60,000/n$ samples per node for MNIST and $50,000/n$ for CIFAR-10.
+We set $\alpha=0.9$ to create heterogeneous distributions, as illustrated in Figure 5 for 16 nodes.
+Evaluation Metrics.
+We measure convergence using the normalized gradient norm computed after each batch.
+Specifically, we average the gradients from all $n$ nodes and normalize by the square root of the total parameter count:
+$\displaystyle\left\|\frac{1}{n}\sum_{i=1}^{n}\mathbf{g}_{i,B_{j}}\right\|_{2}\Big/\sqrt{P},$ (57)
+where $\mathbf{g}_{i,B_{j}}$ is the gradient computed on the $j$-th batch of node $i$, and $P$ is the number of parameters.
+For MNIST with three-layer FCNs, $P=1,863,690$; for CIFAR-10 with ResNet-18, $P=11,173,962$.
+Experimental Setup.
+We test our method on various network topologies including directed exponential and grid graphs, as well as undirected random graphs.
+Training uses learning rate $lr=0.005$ and batch size $B=128$ for MNIST, $B=128$ for CIFAR-10, consistent across all topologies and node configurations.
+Performance is compared against single-node baseline training ($n=1$).
+Figures 2 and 3 demonstrate that Push-Pull achieves linear speedup on both synthetic data and real datasets across directed and undirected topologies.
+Figure 4 shows the validation accuracy comparison for CIFAR-10.
+Figure 5: Comparison of uniform (left) and heterogeneous (right) data distributions across $16$ nodes ($\alpha=0.9$).